### v1

# COBias and Debias: Minimizing Language Model Pairwise Accuracy Bias via Nonlinear Integer Programming

Abstract
For language model classification, would you prefer having only one workable class or having every class working? The latter makes more practical uses. Especially for large language models (LLMs), the fact that they achieve a fair overall accuracy by in-context learning (ICL) obscures a large difference in individual class accuracies. In this work, we uncover and tackle language models’ imbalance in per-class prediction accuracy by reconceptualizing it as the Contextual Oddity Bias (COBias), and we are the first to engage nonlinear integer programming (NIP) to debias it. Briefly, COBias refers to the difference in accuracy by a class A compared to its “odd” class, which holds the majority wrong predictions of class A. With the COBias metric, we reveal that LLMs of varied scales and families exhibit large per-class accuracy differences. Then we propose Debiasing as Nonlinear Integer Programming (DNIP) to correct ICL per-class probabilities for lower bias and higher overall accuracy. Our optimization objective is directly based on the evaluation scores by COBias and accuracy metrics, solved by simulated annealing. Evaluations on three LLMs across seven NLP classification tasks show that DNIP simultaneously achieves significant COBias reduction (-27%) and accuracy improvement (+12%) over the conventional ICL approach, suggesting that modeling pairwise class accuracy differences is a direction in pushing forward more accurate, more reliable LLM predictions.
1 Introduction
We rethink language models’ imbalance in per-class prediction accuracy and reconceptualize it as the Contextual Oddity Bias (COBias). In a nutshell, COBias refers to the difference in accuracy by a class A compared to its “odd” class, which holds the majority wrong predictions of class A. COBias reflects a common language model failure, seen in a variety of NLP benchmark tasks from text classification to question answering. Especially for large language models (LLMs), the fact that they achieve remarkable test accuracy over all class labels by few-shot in-context learning (ICL) [1] can obscure a large difference in individual class accuracies. It is a direct outcome of LLMs’ liking of specific patterns or answers, leading to highly imbalanced predicted class distributions [2, 3]. This per-class accuracy difference can be further attributed to ICL prompt sensitivity or rooted in the black-box pretraining data. To name a few, the causes encompass predilections for common tokens in the pretraining corpora [4], choices of prompt templates [5, 6, 7], demonstrations [8], and prompt orders [9, 10]. Unintended LLM behaviors may also exacerbate the bias, such as user-following sycophancy [11, 12, 13] and prompt toning effects [14, 15, 16, 17]. Despite the causes, COBias renders a language model less effective for lower-accuracy classes and hinders users from trusting the answers, heightening the need for a unified debiasing method at output level.
Importantly, it is not just that a model tends to over-predict a class, but more precisely, the model tends to predict certain classes (not always all others) as it. That is why we need to model the interaction between pairwise class predictions for better bias reduction. An example of an odd class and the per-class accuracy difference are shown by the upper plots in Figure 1. Essentially, mitigating per-class accuracy bias is a critical step in pushing forward LLM inference abilities, as it directly targets improving lower-accuracy classes while improving or at least not hurting the overall accuracy.
In this work, we first put forward COBias and propose the COBias evaluation metric, which measures the averaged pairwise class accuracy difference (Section 2). On three widely used, different-scale LLMs including GPT-2-XL [18], Llama-2-7B and Llama-2-13B [19], we manifest that large COBias consistently exists across seven varied NLP classification tasks.
To reduce COBias and rectify LLM outputs, we directly operate on the per-class probabilities given by the conventional ICL approach. We propose DNIP, the Debiasing as Nonlinear Integer Programming method, to adjust per-class probabilities towards lower COBias and higher overall accuracy, from a combinatorial optimization viewpoint (Section 3). We formulate the debiasing problem as a discrete correction weight selection problem. In particular, DNIP re-weights per-class probability with an optimal set of weights selected based on the evaluation metrics of accuracy and COBias, and thus obtains much less biased predictions with higher accuracy.
On three LLMs across seven NLP evaluation datasets, we show the effectiveness of DNIP. Our key key messages are as follows:
  • We introduce the COBias evaluation metric to assess the pairwise class accuracy bias in language model predictions, and reveal that LLMs of different scales and families consistently exhibit a large COBias score, at an average of 43%.
  • We propose Debiasing as Nonlinear Integer Programming (DNIP) to jointly minimize COBias and maximize accuracy. An efficient simulated annealing algorithm is adopted to solve the mathematical model.
  • Experiments show that DNIP achieves the best of both worlds for lower COBias (avg. $43\%\rightarrow 16\%$) and higher overall accuracy (avg. $52\%\rightarrow 64\%$) over the ICL approach.
2 The Contextual Oddity Bias and the Debiasing Problem
We conceptualize an odd class to be a class at inference time where predictions of another class are often biased towards it. An odd class is relative to the other class. The existence of odd classes manifests the per-class accuracy bias. We term the accuracy difference between a class A and its odd class as the Contextual Oddity Bias (COBias), suggesting a model’s tendency to over-predict an answer that is caught as odd than the correct answer. Notably, an odd answer may reflect surface patterns captured from the task-specific prompts or a common token in the pretraining data; therefore, even on the same task, which answers are odd can vary across different models. That’s why we use “contextual” to mean that the oddity of an answer depends on the model or task. In addition, the prefix “CO” in COBias can be interpreted as the pairwise nature of the bias, suggesting that if there is an odd class, there has to be a class that is biased to it.
The Contextual Oddity Bias Metric. We define the contextual oddity bias metric as follows. Formally, given an $N$-way classification task, let $x_{m}$ denote an input whose label is $y_{m}$. By few-shot prompting, let $\boldsymbol{\boldsymbol{p}_{m}}=(p_{m1},\dots,p_{mN})$ denote the output token probability over $N$ classes $c_{1},\dots,c_{N}$. The ICL prediction $\hat{y}_{m}$ is $\operatorname*{arg\,max}_{j\in\{1,\dots,N\}}\boldsymbol{p}_{m}$. For a given class $c_{i}$, we denote the class it is most biased towards as its oddest class $c_{o}$. Suppose test accuracy for the two is $A_{o}$ and $A_{i}$ respectively, then COBiassingle is computed by:
$\textrm{COBias}_{\textrm{single}}=\displaystyle\sum_{i\neq o}\bigg{|}A_{o}-A_{i}\bigg{|}$ (1)
Class accuracy $A_{i}$ is given by $\frac{1}{|\mathcal{C}_{i}|}\sum_{m\in\mathcal{C}_{i}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\}$, where $\mathcal{C}_{i}$ is the set of indices of class $c_{i}$ examples, and $\mathbbm{1}(\boldsymbol{\cdot})$ returns 1 if the condition inside is satisfied and 0 otherwise.
However, bedsides the oddest class $c_{o}$, there could be the second oddest, the third oddest, or etc. class for a given class. Generalizing the definition, we extend the contextual oddity bias metric as:
$\textrm{COBias}={N\choose 2}^{-1}\displaystyle\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\bigg{|}A_{i}-A_{j}\bigg{|}$ (2)
The contextual oddity bias metric reflects the inter-class imbalance in accuracy. This metric considers each class as the odd class to others, representing the difference between every pair of classes.
The Debiasing Problem. The problem is two-fold - we need to reduce the contextual oddity bias, while maintaining or increasing the overall test accuracy.
3 Improving LLM Outputs by Debiasing as Nonlinear Integer Programming
Figure 1: An overview of the DNIP method. We take the TREC [20] task for illustration; the ICL approach uses 1-shot prompting whose prompt is given by light yellow text boxes. The upper-left plot shows that Loc. is an odd class to the others, holding their majority wrong predictions, especially for Ent.; the upper-right plot demonstrates the stark contrast in per-class accuracy, with 100% for Loc. and 0 for Ent.. The bottom plots show that DNIP greatly reduces pairwise class accuracy differences, especially boosting accuracy for Ent. ($0\rightarrow 66\%$).
In this work, we approach the debiasing problem as a nonlinear integer programming (NIP) problem. The objective is to optimize correction weights to rectify the biased, per-class probabilities, with directly modeling the COBias metric and two accuracy-related metrics. The weights are optimized on a set of labeled examples, and are simply plugged in to inference-time probabilities. An overview of the proposed DNIP method is shown in Figure 1. Our source code will be released.
3.1 The DNIP Objective
Considering $M$ examples with known class labels, we let $\boldsymbol{p}_{m}$ denote the normalized output probability over $N$ class tokens for an example, and $\boldsymbol{p}_{m}=(p_{m1},\dots,p_{mn},\dots,p_{mN})$, $\sum_{n=1}^{N}p_{mn}=1$, $m\in\{1,\dots,M\}$. To debias $\boldsymbol{p}_{m}$, we introduce a discrete $K$-point correction scale $\boldsymbol{\omega}$ ranging from 0 to 1, to re-weight class probability. Let $\boldsymbol{\omega}=(\omega_{1},\dots,\omega_{K})$ denote the parameter for the discrete weights in the correction scale, where $\omega_{k}=k/K$. An example $10$-point scale is given by $\boldsymbol{\omega}=(0.1,\dots,1.0)$. For each class probability in $\boldsymbol{p}_{m}$, it times a correction weight chosen from $\boldsymbol{\omega}$ to obtain a debiased class probability.
Finding optimal correction weights is a combinatorial optimization problem: for $N$-way class probability $\boldsymbol{p}_{m}$, each class selects a weight from the $K$-dimensional correction scale. To find the optimal combination, the brute force strategy is an enumerated search, where the number of weight combinations to explore is $K^{N}$ for an example; for $M$ examples, it results in a computational complexity of $\mathcal{O}(K^{N}\cdot M)$, which can not run in polynomial time, calling for smarter methods.
Instead, we model the debiasing problem as a nonlinear integer program, and search for an optimal set of class correction weights with simulated annealing. To this end, the problem boils down to selecting optimal weight indices, and we define an integer selection variable $\xi=(\xi_{1},\dots,\xi_{N})$ that selects a weight index $k$ for the $n$-th class probability in $\boldsymbol{p}_{m}$, namely:
$\xi_{n}=k\quad\text{if the $n$-th class selects }\omega_{k}.\$ (3)
The DNIP objective function and constraints are formally given by:
$\min\quad$ $Z(\xi)=\frac{1}{M}\displaystyle\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}+\beta{N\choose 2}^{-1}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\bigg{|}A_{i}-A_{j}\bigg{|}-\tau\displaystyle\sum_{j=1}^{N}\textrm{PMI}_{j}$ (4)
s.t. $\hat{y}_{m}=\operatorname*{arg\,max}_{n\in\{1,\dots,N\}}\omega_{\xi_{n}}p_{mn},\qquad m=1,\dots,M$ (5)
$A_{i}=\frac{1}{|\mathcal{C}_{i}|}\sum_{m\in\mathcal{C}_{i}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\},\qquad i=1,\dots,N$ (6)
$\textrm{PMI}_{j}=\textrm{PMI}(\hat{S}^{j},S^{j})=\log\frac{f(\hat{S}^{j},S^{j})}{f(\hat{S}^{j})f(S^{j})},\qquad j=1,\dots,N$ (7)
$\xi_{n}\in\{1,\dots,K\},\qquad n=1,\dots,N$ (8)
where $\beta$ and $\tau$ are parameters that control the bias and the negative PMI terms respectively, $\hat{S}^{j}$ and $S^{j}$ denote the examples of prediction $j$ and true label $j$ respectively [^1] In details, $f(\hat{S}^{j})$ is the ratio between the number of examples with prediction $j$ and the total number of examples, similarly, $f({S}^{j})$ is the ratio between the number of examples with class label $j$ and the total number of examples, $f(\hat{S}^{j},S^{j})$ is the ratio between the number of correct predictions of class $j$ and the total number of examples. Therefore, $\textrm{PMI}_{j}=\log\big{(}\frac{\mathrm{n}(\hat{S}^{j},S^{j})/N}{\mathrm{n}(\hat{S}^{j})/N\cdot\mathrm{n}(S^{j})/N}\big{)}$, where $\mathrm{n}(\cdot)$ is the count. In actual computations, we apply add-$u$ smoothing to smooth out zeros in the numerator and denominator. The value of the smoothing coefficient $u$ is selected along other parameters on a development set.. The debiased class probabilities and prediction are:
$\boldsymbol{p}^{\ast}_{m}$ $=(\omega_{\xi_{1}}p_{m1},\dots,\omega_{\xi_{N}}p_{mN})$ (9)
$y^{\ast}_{m}$ $=\operatorname*{arg\,max}_{n\in\{1,\dots,N\}}\boldsymbol{p}^{\ast}_{m}$ (10)
The intuition is to encourage the adjusted predictions to have lower inter-class bias while improving overall accuracy. As the second term COBias has minimized the inter-class accuracy bias, the first term error rate brings the class prediction closer to the actual class, and the third term negative PMI further enforces the class prediction to be more confident when it is close to the actual class.
3.2 Solving DNIP with Simulated Annealing
Algorithm 1 Optimizing the selection of correction weights with simulated annealing
1:$(\xi,\boldsymbol{p},y,\omega)$: weight selection variable, class probabilities and labels, the weight scale
2:$(\xi,\{y^{\ast}_{m}\}_{1}^{M})$
3:$T\leftarrow T_{max}$, $\xi^{\ast}\leftarrow\xi\leftarrow$ INIT()
4:while $T\geq T_{min}$ do
5:     while inner-loop criterion is not satisfied do
6:         $\xi_{new}\leftarrow$ PERTURB($\xi$$\triangleright$ Sample a new $\xi$
7:         $\Delta z\leftarrow$ $z(\xi_{new})-z(\xi)$
8:         if $\Delta z\leq 0$ then
9:              $\xi\leftarrow$ $\xi_{new}$
10:              if $z(\xi_{new})<z(\xi^{\ast})$ then
11:                  $\xi^{\ast}\leftarrow$ $\xi_{new}$
12:         else if random($0,1$) $<$ exp($\Delta z/T$then
13:              $\xi\leftarrow$ $\xi_{new}$triangleright Accept a worse $\xi$
14:     $T\leftarrow$ $\alpha T$
15:$y^{\ast}\leftarrow$ INFER($\xi^{\ast},\boldsymbol{p},\omega$)
16:return $\xi^{\ast},y^{\ast}$
We solve DNIP by simulated annealing [21, 22]. Simulated annealing (SA) is a Metropolis-Hastings sampling algorithm, which is widely adopted to solve optimization problems with its versatility in dealing with highly nonlinear models, noisy data, and many constraints [23]. The DNIP mathematical model is such a case, pointing to the need of SA. We especially take advantages of SA algorithms’ searching ability - escaping from local optima and finding global optima after reaching local optima - to find optimal correction weights. In details, the SA strategy consists of searching the solution space starting from a randomly initialized $\xi$ and then generating a new one by perturbing it. The new solution is either accepted or rejected by the Metropolis criterion after evaluating the objectives.
Algorithm 1 describes the searching for optimal $\xi$ using SA. The outer loop performs the cooling/annealing process. In this work, a geometric decay with $\alpha=0.95$ and initial temperature of 200,000 is set as cooling schedule. The inner loop simulates the thermal equilibrium reaching process at a temperature, and its criterion ensures that either the number of generated or accepted solutions at a temperature is above a threshold [24]. We sample a new $\xi$ from the neighborhood by randomly substitute a selection in $\xi$. The computational complexity for the SA algorithm is $\mathcal{O}(NK)$. Please refer to the Appendix Section A for derivations.
4 Experiments
4.1 Experimental Setup
Evaluation Tasks. The proposed method is evaluated on a diverse set of binary and multi-class NLP classification tasks, across general domains and biomedical domains. The five general-domain evaluation datasets are 4-way news topic classification AGNews [25], 14-way ontology classification DBpedia [25], 5-way sentiment classification SST-5 [26], 6-way retrieval question classification TREC [27, 20], binary entailment recognition RTE [28]; the two domain-specific datasets from BLURB [29] are 5-way biomedical relation extraction DDI [30], and 3-way biomedical question answering PubMedQA [31]. The selected biomedical tasks are of practical uses. For example, detecting DDI is useful in preventing adverse effects from drug combinations.
Each dataset comes with a training and a test set. We make use of the default training set for DNIP optimization, except for AGNews, DBpedia, and DDI, for which we randomly select 10,000 training examples since the datasets are quite large. We then split the training examples into 95%, 5% optimization and development sets. We evaluate on the default test set except for AGNews, for which we evaluate on 5,000 randomly selected test examples. Evaluation metrics are COBias and accuracy.
Models and Experimental Configuration. We evaluate three open-source LLMs across varied scales and model families, including GPT-2-XL (1.5B parameters), Llama-2-7B (7B parameters), and Llama-2-13B (13B parameters). For ICL, we use 1-shot prompting, and obtain the output per-class softmax probabilities. To standardize calculations, we normalize the softmax probabilities over all classes. For DNIP, we tune parameters of $\beta,\tau,K,u$ on the development set. To account for variance caused by different demonstrations, for each model and dataset, we perform three runs of prompting with different 1-shot demonstrations and obtain three sets of initial probabilities. For both ICL and DNIP, we report the average test accuracy and COBias over the three runs. The demonstration is randomly selected from training examples. All prompting is done on an NVIDIA A100 GPU. The simulated annealing algorithm executes on CPU (execution time depends on the optimization set size, 3 to 30 minutes on AMD EPYC 7742 CPU or slightly longer on local laptop CPU).
Model Eval. Metric AGNews DBpedia SST-5 TREC RTE DDI PubMedaQA Avg.
1-shot ICL
GPT-2-XL
Acc
COBias
52.15.4
35.511.5
31.89.9
40.03.6
34.913.7
48.75.4
27.410.5
45.68.7
55.41.9
82.424.5
14.54.4
40.75.9
55.20.0
59.412.6
38.8
50.3
Llama-2-7B
Acc
COBias
86.42.5
14.06.5
88.92.0
13.52.1
42.111.1
55.61.5
66.76.6
33.210.0
66.34.3
61.610.5
6.70.4
41.41.7
40.36.7
40.916.1
56.8
37.2
Llama-2-13B
Acc
COBias
79.97.0
28.316.1
88.61.7
16.23.6
44.94.3
53.15.0
68.510.8
35.96.5
71.52.2
43.47.0
7.20.9
45.65.9
55.12.9
61.21.9
59.4
40.5
DNIP
GPT-2-XL
Acc
COBias
68.51.0
1.40.5
69.99.1
24.18.3
44.52.20
26.02.5
46.312.7
27.27.2
50.82.1
7.15.0
43.914.9
17.07.1
57.11.3
29.825.0
54.4
18.9
Llama-2-7B
Acc
COBias
86.70.4
1.30.1
92.90.4
7.70.6
50.62.7
28.021.6
68.11.0
15.91.6
73.92.3
1.91.8
44.53.8
11.63.0
62.78.3
35.422.8
68.5
14.5
Llama-2-13B
Acc
COBias
87.90.7
6.30.6
93.40.6
7.70.6
48.31.9
18.710.1
77.12.0
14.21.3
74.20.7
4.33.4
40.46.0
7.53.2
63.114.0
41.129.6
69.2
14.3
Table 1: Performance comparisons between the 1-shot ICL approach and DNIP. In each cell, we report the average test accuracy/COBias (%) and standard deviation over three different prompts. DNIP greatly improves test accuracy and significantly lowers test COBias upon the ICL approach.
4.2 Main Results
The average test accuracy (in black) and COBias (in blue) with standard deviation over three runs are shown in Table 1. We find that test COBias by ICL is large in all three LLMs, with GPT-2-XL having the largest average COBias of 50.3%, showing that LLMs of different sizes and families exhibit large per-class accuracy differences. With DNIP, the average test COBias over three models reduces from 43% to 16%, and the test accuracy increases from 52% to 64%. Specifically, DNIP achieves an 16%, 12%, and 10% absolute improvement in test accuracy for GPT-2-XL, Llama-2-7B, and Llama-2-13B respectively; meanwhile, DNIP significantly reduces test COBias by an absolute 29%, 23%, and 26% for the three models, over ICL results. As for detailed results, for GPT-2-XL where the ICL test accuracy is below 30% (DDI), DNIP boosts test accuracy by 29% while lowering test COBias from 41% to 17%; for Llama-2 7B and 13B where the ICL test accuracy is already in higher 80s, e.g., DBpedia, DNIP can further improve the accuracy to 90s while reducing COBias. In addition, we observe a very large ICL test COBias at 82.4% for GPT-2-XL on RTE. After applying DNIP, the COBias greatly reduces to 7.1%. The 82.4% COBias on RTE suggests that the conventional ICL predictions mostly fall in a single class, failing to predict the other class. In fact, The RTE task seems too challenging that GPT-2-XL’s output probability of class True is much higher than that of class False. Though re-weighting can lower both probabilities, it may not gain a corrected prediction. Still, the over 75% absolute reduction in COBias demonstrates that DNIP in the worst case can balance severely skewed ICL predictions without hurting the overall accuracy much. On two biomedical tasks, DNIP brings an average absolute increase of 34% and 11% in accuracy, and an average absolute reduction of 31% and 13% in COBias, showing its effectiveness for domain-specific tasks.
These results show that DNIP strongly mitigates COBias and greatly improves the accuracy of lower-performing classes, while boosting overall accuracy. Moreover, relatively large LLMs can suffer even higher COBias than smaller models on biomedical tasks, suggesting that the per-class accuracy bias does not go away as models scale, and debiasing is essentially needed to improve inference for larger LLMs.
4.3 DNIP Balances Class Accuracies across Varied Classification Tasks, as many as 14 Classes
Figure 2: Comparisons of pairwise class accuracy differences on test sets before and after applying DNIP; the pinker the higher difference, the greener the lower difference; average accuracy over three runs are used. Heatmaps after debiasing show nearly 0 differences for many class pairs, demonstrating DNIP’s effectiveness for COBias reduction.
We visualize pairwise class accuracy differences before and after applying DNIP on Llama-2-13B by the heatmaps in Figure 2, which show that DNIP is straightforwardly effective in reducing COBias. In each heatmap, the value of row $i$, column $j$ cell represents the absolute difference between the test accuracies of two classes $i,j$ , i.e., $|A_{i}-A_{j}|$. We observe that DNIP brings the accuracy differences down to close to zero for 4 out of 7 datasets, including AGNews, DDI, DBpedia, and RTE; and the most relative reduction is seen on TREC. These results show that DNIP consistently reduces prediction bias across different NLP classification tasks.
Figure 3: Test confusion matrix for DBpedia before and after applying DNIP. DNIP greatly reduces wrong predictions for Nat., which is the lowest-accuracy class before debiasing.
Relieving Odd Classes for 14-Way Ontology Classification. We further investigate the 14-way classification task DBpedia. Before debiasing, the Nat. class exhibits the largest class accuracy difference compared to other classes, suggesting that Nat. is either the highest-accuracy class manifesting a strong prediction bias towards Nat., or the lowest-accuracy class where predictions of true Nat. examples are biased to one or more odd classes. We plot the confusion matrix of test results of 1 run out of 3 runs in Figure 3 and find that Nat. has the lowest accuracy, and its predictions are biased to two odd classes, Trans. and Bldg., with Trans. being the oddest to Nat.. DNIP greatly reduces false negatives for Nat., i.e., reduces wrong Trans. and Bldg. predictions for Nat.. Besides these visualizations, on an average over three runs, the test accuracy for Nat. improves from 44% to 89%, and that for Trans. also increases from 78% to 87%, demonstrating DNIP’s effectiveness in reducing pairwise class accuracy bias for classification tasks of as many as 14 classes.
Furthermore, we see more balanced class accuracies with DNIP for smaller LLMs of Llama-2-7B and GPT-2-XL, in Appendix Section B.
4.4 DNIP Objective Ablation
We present ablation analysis on objective functions to show that the integration of accuracy-related objectives and the COBias objective are indispensable in achieving the best of both worlds. Recall three parts in our objective function $z$: the error rate term $z_{1}=\frac{1}{M}\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}$; the COBias term $z_{2}={N\choose 2}^{-1}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}|A_{i}-A_{j}|$; and the PMI term $z_{3}=\sum_{j=1}^{N}\textrm{PMI}_{j}$; the accuracy-related terms are $z_{1}$ and $z_{3}$. We set up 7 different objective functions to perform DNIP: $z_{1},z_{2},z_{3},z_{1}+\beta z_{2},z_{1}-\tau z_{3},\beta z_{2}-\tau z_{3},z_{1}+\beta z_{2}-\tau z_{3}$, and compare their effectiveness on Llama-2-13B in Figure 4. Ablations on GPT-2-XL and Llama-2-7B show similar findings (Appendix Section C). To make fair comparisons, we fix $\beta=2.7,\tau=0.2$ and weight scale $K=30$, which may not result in the same scores reported in Table 1.
Figure 4: Ablation analysis of the objective function, exemplified with Llama-2-13B; test scores averaged over 7 datasets are used for the rightmost plot.
Among all ablations, $z_{1}+\beta z_{2}-\tau z_{3}$ achieves a balance point between accuracy and COBias. For the criteria, lower COBias is our top priority; when COBias scores are similarly low, we prefer the one that achieves higher accuracy. Compared to $z_{1}+\beta z_{2}-\tau z_{3}$$z_{1}$: (1) Using $z_{1}$ or $z_{3}$ solely emphasizes improving accuracy; though obtaining higher or similar accuracy, they suffer a higher COBias. (2) Using $z_{1}-\tau z_{3}$ also obtains a higher/similar accuracy but a higher COBias. (3) Using $z_{2}$ solely focuses on mitigating COBias and is straightforwardly effective in reducing COBias. However, the shortcoming is obvious: a lower accuracy. (4) Using $z_{1}+\beta z_{2}$ or $\beta z_{2}-\tau z_{3}$ is promising with a similar/lower accuracy and a slightly lower COBias. We note that accuracy-related objectives alone increase accuracy by over 16% ($75\%-59\%$), which is greater than using the objective combining COBias, suggesting the NIP can be a direction for those who look for accuracy improvement.
4.5 Performance Improvements vs. Number of Optimization Examples
We evaluate DNIP with varying optimization examples on Llama-2-13B. We find that DNIP requires only a small optimization set of 10 examples for a fair improvement in both COBias and accuracy, and it exhibits a further, emergent COBias reduction at several thousands optimization examples.
Figure 5: Relative test COBias reduction; higher is better.
Test COBias reduction reaches 9% to 72% at 10 optimization examples, and continues to grow with more optimization examples. Figure 5 shows the relative test COBias reduction over ICL results. At 10 examples, DNIP can achieve a relative COBias reduction ranging from 9% to 72%. We observe an overall trend of COBias reduction except for DDI and PubMedQA, who have a spike at 100 examples, but DDI rises again after 500 examples while PubMedQA levels off (PubMedQA full optimization size: 950). An explanation is that DNIP prioritizes COBias reduction more than accuracy increase at fewer examples; at more examples, the accuracy-related objectives show more influence. It also indicates that domain-specific datasets are more sensitive to the trade-off. We further notice a surge in reductions from 1,000 to full for DDI, SST-5, and DBpedia (full optimization sizes: 9,500, 8,116, 9,500), suggesting that around 8,000 examples are needed for a further emergent reduction - PubMedQA may see further COBias reduction if more optimization examples are available.
Figure 6: Relative test accuracy increase; higher is better.
Test accuracy is improved with only 10 optimization examples, and plateaus at around 500 examples. Figure 6 shows that, at 10 examples, the relative increase in accuracy ranges from 0.2% to 11% except DDI, which is 160%. At 100 examples, the increase ranges from 3% to 208% (DDI, $7.2\%\rightarrow 18.7\%$). The increase continues to grow and stabilizes at around 500 examples. At 1,000 examples, test accuracy for DDI rises to 52%, which is more than 6 times of the ICL accuracy, showing that DNIP is effective for this challenging biomedical NLP task. Moreover, we observe for RTE a declining trend of accuracy increase, which matches its growing trend of COBias reduction, reflecting the trade-off between the two objectives. We also note a slight decrease in accuracy improvement for SST-5 and DDI at full size, which matches their sudden increase in COBias reduction, verifying that DNIP has emergent COBias reduction abilities with thousands of optimization examples. Overall, DNIP improves both accuracy and COBias at 10 examples, and achieves acceptable COBias reduction and accuracy improvement at 100 examples, with test COBias reduced by 57% and test accuracy improved. For detailed scores, please refer to Appendix D.
5 Related Work
LLMs may fail in various ways. Analysis towards understanding LLM failure modes has gained remarkable insights on how they fail and how to identify/avoid failures. For example, understanding hallucinations in LLMs facilitates fact-checking evaluation metrics for LLMs.[32, 33, 34]. Jones and Steinhardt [14] reveals failure modes of LLMs which resemble cognitive biases and measures the failures on GPT-3.5 and Codex. Lu et al. [9] identify ICL order sensitivity that result in LLM prompting failures. Turpin et al. [10] shows that LLM reasoning abilities are affected by adding biasing features to the input. Along this line, we tackle a common LLM failure of the per-class accuracy difference, via nonlinear integer programming optimization on the evaluation metric scores.
6 Discussion
Using Chain-of-Thought Prompting to Improve LLM Predictions. Some may argue that a simple remedy for biased predictions is Chain-of-Thought (CoT) prompting, which elicit more reasoning in LLMs. However, Turpin et al. [10] points out that CoT explanations can be unfaithful, as they can be biased towards class A when prompts always put letter A as the answer, causing 36% accuracy drops on BBH [35]. We add that prompt-based models may fall for pattern matching instead of recalling learned factual knowledge [2]. Therefore, CoT may not help due to unfaithful explanations.
Using Calibration Methods to Mitigate ICL Biases. Previous calibration methods reduce biases caused by prompting. For example, Zhao et al. [4], Zhou et al. [36] use content-free or content-based prompts to measure an offset and calibrate class probabilities by removing it. This may be good for mitigating bias towards a single class, but it does not consider the interaction between class predictions, e.g., pairwise class accuracy bias studied in this work. Another hazard is using prompts to compute the offset, which may cascade biases brought by a certain or several prompt templates, and biases inherited from pretraining may not be reflected. Another line of works does not rely on further prompting. Han et al. [3] fits prototypical clusters on class probabilities and updates predictions by a cluster-label mapping. We share the same spirit of correcting biased predictions without more prompting, but we differ in that our correction weights are explicitly optimized on COBias and accuracy metrics, while the cluster-based decision boundary is implicitly learned.
Evaluating Closed LLMs. While closed LLMs such as ChatGPT and GPT-4 [37] are interesting to evaluate, they do not return logits/softmax probabilities over the entire vocabulary at an output token. For example, ChatGPT returns log probabilities of up to 20 most likely output tokens, and may not cover the probability for every class that is needed in our class prediction debiasing.
7 Conclusion
This work tackles the imbalance in per-class prediction accuracy of language models. We introduce COBias to assess pairwise class accuracy differences, and reveal that LLMs of varied scales and families exhibit large COBias at inference. To minimize COBias and push for higher accuracy, we view debiasing as a combinatorial optimization problem and propose DNIP, which optimizes class correction weights by jointly minimizing evaluation scores by COBias and maximizing accuracy. Empirical evaluations on three LLMs across seven multi-class NLP classification tasks demonstrate that DNIP simultaneously achieves lower test COBias and higher test accuracy over conventional ICL. We advocate that reducing COBias is a direction to gain more accurate, reliable LLM predictions.
Limitations and Future Work. The limitation is that we only evaluate NLP classification tasks. We leave evaluations on open-ended generations and other modalities for future work. Indeed, DNIP is fruitful for debiasing imbalanced class predictions, which can easily extend to vision/multimodal LMs, as long as there is a need for mitigating per-class accuracy differences. This simple inference-time plug-in solution is there for further boosting LM performances.
Appendix A Derivation of SA Algorithm Complexity
[…]
Appendix B Additional Heatmaps for Pairwise Accuracy Differences
Figure 7 and 8 demonstrate that more balanced class accuracies are achieved with DNIP for smaller LLMs of Llama-2-7B and GPT-2-XL.
Figure 7: Comparisons of pairwise test class accuracy differences before and after applying DNIP on GPT-2-XL; the pinker the higher accuracy difference, the greener the lower accuracy difference; average accuracy over three runs are used.
Figure 8: Comparisons of pairwise test class accuracy differences before and after applying DNIP on Llama-2-7B; the pinker the higher accuracy difference, the greener the lower accuracy difference; average accuracy over three runs are used.
Appendix C Additional Ablation Analysis
We show ablation analysis on GPT-2-XL and Llama-2-7B in Figure 9 and 10.
Figure 9: Ablations on objective function, GPT-2-XL, demonstrating that $z_{1}+\beta z_{2}-\tau z_{3}$ achieves a balance point between accuracy and COBias.
Figure 10: Ablations on objective function, Llama-2-7B, demonstrating that $z_{1}+\beta z_{2}-\tau z_{3}$ achieves a balance point between accuracy and COBias.
Appendix D Full Table with Varying Number of Optimization Examples for DNIP
The full results for all three models across varying optimization set sizes are shown in Table 2 (on the next page).
Opt. set Eval. Metric AGNews DBpedia SST-5 TREC RTE DDI PubMedaQA
GPT-2-XL
0
Acc
COBias
52.15.4
35.511.5
31.89.9
40.03.6
34.913.7
48.75.4
27.410.5
45.68.7
55.41.9
82.424.5
14.54.4
40.75.9
55.20.0
59.412.6
10$\ast$
Acc
COBias
72.82.1
17.21.7
40.65.2
28.06.6
40.33.4
40.11.8
51.55.7
39.71.0
50.30.8
6.57.6
23.51.0
40.63.5
53.70.6
41.80.0
100
Acc
COBias
70.05.0
14.81.3
41.429.9
27.88.9
44.31.3
30.83.4
49.010.1
31.15.9
49.82.5
5.34.9
42.25.7
17.17.1
46.710.8
29.216.6
500
Acc
COBias
68.21.2
2.71.8
44.533.2
25.79.6
43.41.5
29.83.7
49.98.5
31.46.2
50.52.4
5.23.9
45.69.2
24.514.3
49.911.4
17.715.8
1,000
Acc
COBias
69.61.0
3.01.2
42.032.2
27.611.0
42.82.3
28.25.9
50.99.0
24.55.4
50.32.2
5.23.3
48.34.9
22.92.5
same as
full
Full
Acc
COBias
68.51.0
1.40.5
69.99.1
24.18.3
44.52.20
26.02.5
46.312.7
27.27.2
50.82.1
7.15.0
43.914.9
17.07.1
57.11.3
29.825.0
Llama-2-7B
0
Acc
COBias
86.42.5
14.06.5
88.92.0
13.52.1
42.111.1
55.61.5
66.76.6
33.210.0
66.34.3
61.610.5
6.70.4
41.41.7
40.36.7
40.916.1
10$\ast$
Acc
COBias
86.42.5
14.06.5
89.91.4
12.52.1
51.40.9
36.23.6
70.10.9
22.45.4
74.92.1
4.85.1
31.721.0
26.78.2
44.50.6
28.73.9
100
Acc
COBias
88.40.4
5.80.6
91.80.7
9.71.0
50.91.5
34.312.7
70.11.0
16.72.4
73.62.6
2.90.7
44.92.5
21.06.6
62.64.5
35.418.1
500
Acc
COBias
86.80.9
2.80.3
92.10.6
8.61.6
50.81.7
35.812.3
69.71.2
15.22.1
74.32.9
3.11.4
69.31.7
34.50.3
63.57.5
37.320.4
1,000
Acc
COBias
86.80.3
1.90.5
92.50.2
8.00.4
51.02.1
30.920.4
69.50.8
15.71.2
74.02.5
2.61.2
56.57.9
29.53.4
same as
full
Full
Acc
COBias
86.70.4
1.30.1
92.90.4
7.70.6
50.62.7
28.021.6
68.11.0
15.91.6
73.92.3
1.91.8
44.53.8
11.63.0
62.78.3
35.422.8
Llama-2-13B
0
Acc
COBias
79.97.0
28.316.1
88.61.7
16.23.6
44.94.3
53.15.0
68.510.8
35.96.5
71.52.2
43.47.0
7.20.9
45.65.9
55.12.9
61.21.9
10$\ast$
Acc
COBias
86.01.9/
14.33.5
88.81.6/
14.72.5
49.20.8/
41.78.1
75.83.2
29.64.5
74.82.3
12.25.7
18.712.6
30.07.8
59.66.2
35.37.8
100
Acc
COBias
87.80.2/
8.30.5
91.80.2/
10.10.7
47.42.0/
30.44.4
78.32.8
18.52.6
75.51.0
10.46.9
22.22.0
17.34.4
56.910.4
25.816.3
500
Acc
COBias
87.11.5
5.33.1
92.70.0/
9.10.3
49.50.7/
28.03.2
78.72.1
16.31.0
73.61.7
4.03.4
48.227.4
28.011.5
62.215.8
40.429.4
1,000
Acc
COBias
87.70.7/
6.71.5
92.80.1/
9.20.2
49.80.7/
28.82.0
77.32.0
14.21.8
73.91.5
3.64.3
52.010.5
25.09.7
same as
full
Full
Acc
COBias
87.90.7
6.30.6
93.40.6
7.70.6
48.31.9
18.710.1
77.12.0
14.21.3
74.20.7
4.33.4
40.46.0
7.53.2
63.114.0
41.129.6
Table 2: Comparisons of DNIP results on varying optimization set sizes; min. optimization set size is 10 for all datasets except for DBpedia, which we use 15 to cover its 14 classes. For full optimization set sizes, AGNews: 9,500, DBpedia: 9,500, SST-5: 8,116, TREC: 5,179, RTE: 2,365, DDI: 9,500, PubMedQA: 950.

### v3

@@
-COBias and Debias: Minimizing Language Model Pairwise Accuracy Bias via Nonlinear Integer Programming
+Balancing Class Accuracies for Language Models in Inference Time via Nonlinear Integer Programming
 Abstract
-For language model classification, would you prefer having only one workable class or having every class working?
-The latter makes more practical uses.
-Especially for large language models (LLMs), the fact that they achieve a fair overall accuracy by in-context learning (ICL) obscures a large difference in individual class accuracies.
-In this work, we uncover and tackle language models’ imbalance in per-class prediction accuracy by reconceptualizing it as the Contextual Oddity Bias (COBias), and we are the first to engage nonlinear integer programming (NIP) to debias it.
-Briefly, COBias refers to the difference in accuracy by a class A compared to its “odd” class, which holds the majority wrong predictions of class A.
-With the COBias metric, we reveal that LLMs of varied scales and families exhibit large per-class accuracy differences.
-Then we propose Debiasing as Nonlinear Integer Programming (DNIP) to correct ICL per-class probabilities for lower bias and higher overall accuracy.
-Our optimization objective is directly based on the evaluation scores by COBias and accuracy metrics, solved by simulated annealing.
-Evaluations on three LLMs across seven NLP classification tasks show that DNIP simultaneously achieves significant COBias reduction (-27%) and accuracy improvement (+12%) over the conventional ICL approach, suggesting that modeling pairwise class accuracy differences is a direction in pushing forward more accurate, more reliable LLM predictions.
+Large language models are good knowledge bases but struggle to perform equally well for all classes in simple text classification tasks.
+This paper investigates a fundamental inference-time problem in language models: imbalanced class accuracies.
+We find what’s underneath the issue is a tendency to over-predict some classes while under-predicting some others, for both base LLMs and more sophisticatedly trained, larger chat LLMs.
+This class accuracy imbalance is difficult to solve from the root via “better” pre-training or fine-tuning strategies, but we show it can be effectively mitigated via inference-time optimization.
+To this end, we conceptualize and quantify the over- and under-prediction issue as the Contextual Oddity Bias (COBias), and propose the Debiasing as Nonlinear Integer Programming (DNIP) model to correct in-context learned class probabilities based on minimizing COBias and maximizing overall accuracy, without LLM parameter update.
+Considering that the DNIP model implicitly contains non-differentiable elements, we therefore use the simulated annealing algorithm to solve.
+We perform extensive evaluations on three LLMs across seven NLP classification tasks, in different prompting settings.
+Results show that DNIP simultaneously achieves significant COBias reduction (-27%) and accuracy improvement (+12%) over the conventional ICL approach, suggesting that inference-time mitigation of class accuracy imbalance is a promising direction to push forward LLM performances.
 1 Introduction
-We rethink language models’ imbalance in per-class prediction accuracy and reconceptualize it as the Contextual Oddity Bias (COBias).
-In a nutshell, COBias refers to the difference in accuracy by a class A compared to its “odd” class, which holds the majority wrong predictions of class A.
-COBias reflects a common language model failure, seen in a variety of NLP benchmark tasks from text classification to question answering.
-Especially for large language models (LLMs), the fact that they achieve remarkable test accuracy over all class labels by few-shot in-context learning (ICL) [1] can obscure a large difference in individual class accuracies.
-It is a direct outcome of LLMs’ liking of specific patterns or answers, leading to highly imbalanced predicted class distributions [2, 3].
-This per-class accuracy difference can be further attributed to ICL prompt sensitivity or rooted in the black-box pretraining data.
-To name a few, the causes encompass predilections for common tokens in the pretraining corpora [4], choices of prompt templates [5, 6, 7], demonstrations [8], and prompt orders [9, 10].
-Unintended LLM behaviors may also exacerbate the bias, such as user-following sycophancy [11, 12, 13] and prompt toning effects [14, 15, 16, 17].
-Despite the causes, COBias renders a language model less effective for lower-accuracy classes and hinders users from trusting the answers, heightening the need for a unified debiasing method at output level.
-Importantly, it is not just that a model tends to over-predict a class, but more precisely, the model tends to predict certain classes (not always all others) as it.
-That is why we need to model the interaction between pairwise class predictions for better bias reduction.
-An example of an odd class and the per-class accuracy difference are shown by the upper plots in Figure 1.
-Essentially, mitigating per-class accuracy bias is a critical step in pushing forward LLM inference abilities, as it directly targets improving lower-accuracy classes while improving or at least not hurting the overall accuracy.
-In this work, we first put forward COBias and propose the COBias evaluation metric, which measures the averaged pairwise class accuracy difference (Section 2).
-On three widely used, different-scale LLMs including GPT-2-XL [18], Llama-2-7B and Llama-2-13B [19], we manifest that large COBias consistently exists across seven varied NLP classification tasks.
-To reduce COBias and rectify LLM outputs, we directly operate on the per-class probabilities given by the conventional ICL approach.
-We propose DNIP, the Debiasing as Nonlinear Integer Programming method, to adjust per-class probabilities towards lower COBias and higher overall accuracy, from a combinatorial optimization viewpoint (Section 3).
+The fact that large language models (LLMs) achieve remarkable overall test accuracy by few-shot in-context learning (ICL) [1] can obscure a large difference in individual class accuracies.
+Mitigating class accuracy imbalance is under-explored, but important to facilitate a smoother user experience.
+Not mitigating such imbalances may lead to serious damages in some cases.
+For example, a patient can trust an LLM about drug-drug interaction type predictions without knowing that the model over-predicts “no interaction” and has low accuracies for other categories of interactions.
+If the patient trusts “no interaction” when the ground truth is an actual drug interaction type, taking the drugs together can put the patient’s life at risk.
+Therefore, it is high time we took actions to reduce class accuracy imbalances, better without hurting overall accuracy.
+It is difficult to disentangle all the factors that cause class accuracy imbalance, including biases incurred in pre-training, supervised fine-tuning, RLHF, prompting specifics, etc.
+Therefore, it is challenging to comprehensively resolve the root causes through pre-training or fine-tuning methods, or approaches like data selection or data augmentation.
+To overcome these difficulties, we conceptualize a post-hoc bias metric, the Contextual Oddity Bias (COBias; Section 2), to quantify the class accuracy imbalance based on inference-time over-predictions and under-predictions.
+To reduce COBias and rectify LLM outputs, we directly operate on the class probabilities obtained from in-context learning (ICL).
+We propose DNIP, the Debiasing as Nonlinear Integer Programming model, to adjust class probabilities towards lower COBias and higher overall accuracy, from a combinatorial optimization viewpoint (Section 3).
 We formulate the debiasing problem as a discrete correction weight selection problem.
-In particular, DNIP re-weights per-class probability with an optimal set of weights selected based on the evaluation metrics of accuracy and COBias, and thus obtains much less biased predictions with higher accuracy.
-On three LLMs across seven NLP evaluation datasets, we show the effectiveness of DNIP. Our key key messages are as follows:
-  • We introduce the COBias evaluation metric to assess the pairwise class accuracy bias in language model predictions, and reveal that LLMs of different scales and families consistently exhibit a large COBias score, at an average of 43%.
+In particular, DNIP re-weights class probability with an optimal set of weights selected based on both metrics of accuracy and COBias, and thus obtains much less biased predictions with higher accuracy.
+On three widely used different-scale LLMs across seven NLP evaluation datasets, and in different ICL settings, we show the effectiveness of DNIP. Our key messages are as follows:
+  • We introduce the COBias evaluation metric to assess the pairwise class accuracy bias in language model predictions, and reveal that LLMs of different scales and families consistently exhibit a large COBias score.
   • We propose Debiasing as Nonlinear Integer Programming (DNIP) to jointly minimize COBias and maximize accuracy.
 An efficient simulated annealing algorithm is adopted to solve the mathematical model.
-  • Experiments show that DNIP achieves the best of both worlds for lower COBias (avg. $43\%\rightarrow 16\%$) and higher overall accuracy (avg. $52\%\rightarrow 64\%$) over the ICL approach.
-2 The Contextual Oddity Bias and the Debiasing Problem
-We conceptualize an odd class to be a class at inference time where predictions of another class are often biased towards it.
-An odd class is relative to the other class.
-The existence of odd classes manifests the per-class accuracy bias.
-We term the accuracy difference between a class A and its odd class as the Contextual Oddity Bias (COBias), suggesting a model’s tendency to over-predict an answer that is caught as odd than the correct answer.
-Notably, an odd answer may reflect surface patterns captured from the task-specific prompts or a common token in the pretraining data; therefore, even on the same task, which answers are odd can vary across different models.
+  • Experiments show that DNIP achieves the best of both worlds of lower COBias and higher overall accuracy.
+For example, on Llama-2-13B, DNIP significantly reduces COBias (avg.
+$40.5\%\rightarrow 14.3\%$) and improves overall accuracy (avg.
+$59.4\%\rightarrow 69.2\%$) over the 1-shot ICL results, and it reduces COBias even further over the 5-shot ICL results (avg.
+$39.2\%\rightarrow 10.3\%$).
+2 Over-prediction, Under-prediction, and the Contextual Oddity Bias
+2.1 Over-prediction and Under-prediction in Classification
+[…]
+In the era of Large Language Models (LLMs), over-prediction and under-prediction still persist in LLMs’ predictions, though one may not simply attribute such issues to the black-box training data or the model itself.
+The first obstacle is that the pretraining data’s source distribution for a subgroup of data may not be accessible, making it difficult to directly minimize differences between the source distribution and a downstream target distribution.
+What we normally have is the pretrained model’s output distribution for downstream classes.
+Secondly, the process of prompting an LLM, especially when the prompt is provided with in-context learned demonstrative examples, also brings uncertainty to predictions.
+Ablating the potential biases in prompting is non-trivial.
+More inconveniently, given the relatively large model sizes, it is computationally intensive, expensive, or even infeasible to tweak the LLM training pipeline or conduct LLM fine-tuning to fit debiasing training objectives.
+To see how over-predictions and under-predictions hurt LLM prediction capabilities, we show in Table 2 an example of observed over-predictions and under-predictions in Llama-2-13B.
+We prompt the LLM on the AGNews news classification test set in a one-shot manner, with a single demonstrative example randomly selected from the training set.
+Llama-2-13B greatly over-predicts Business while under-predicting Technology, lowering the LLM’s capability for accurately predicting the Tech class (Tech’s accuracy: 19%).
+In general, there are 2129 predictions in Business but the number of ground-truth instances in Business is 1204; there are only 246 correct Tech predictions but 1240 ground-truth Tech instances.
+Over-/Under-prediction happens in more downstream tasks and larger LLMs (e.g., ChatGPT), as additionally show by examples in Appendix A.
+These surface observations tell us how an LLM could fail in predicting certain classes, suggesting that there is still space for improvements for LLMs to predict every class with equally good accuracy.
+The fact that this problem is difficult to completely eliminate from the source of data, together with the inconveniences in training/fine-tuning debiased LLMs, call for a more actionable inference-time bias metric quantifying the over-/under-predictions and an accompanying debiasing method.
+Pred.
+World Sports Business Tech All Acc.
+World 1093 64 126 3 1286 0.85
+Sports 9 1247 14 0 1270 0.98
+True Business 25 4 1167 8 1204 0.97
+Tech 156 27 822 235 1240 0.19
+All 1283 1342 2129 246 5000 -
+Prec. 0.85 0.93 0.55 0.96 - -
+Table 1: Over-prediction in class Business and under-prediction in class Tech in Llama-2-13B, shown by the test confusion matrix for AGNews.
+Pred.
+World Sports Business Tech All Acc.
+World 1100 50 67 69 1286 0.86
+Sports 16 1240 11 3 1270 0.98
+True Business 44 2 1028 130 1204 0.85
+Tech 55 7 138 1040 1240 0.84
+All 1215 1299 1244 1242 5000 -
+Prec. 0.91 0.95 0.83 0.84 - -
+Table 2: Significantly reduced over-predictions in class Business and improved test accuracy of the under-predicted class Tech for AGNews after DNIP.
+2.2 Conceptualizing Over-prediction and Under-prediction into the Contextual Oddity Bias
+In Table 2, we observe that Llama-2-13B greatly over-predicts Business while under-predicting Technology (Tech).
+Underneath the surface over-/under-prediction issue, we can see the most significant contributor is the 822 ground-truth instances of Technology, which are wrongly predicted as Business, leading to the model’s low accuracy (19%) in Technology [^1] It simultaneously leads to a fairly low precision score of 55% in Business.
+In fact, the low precision score in Business and the low recall/accuracy score in Technology co-reflect each other.
+In other words, class accuracies are essential indicators for class performances.
+All current evaluation metrics are computed by Python scikit-learn.
+Note that by scikit-learn, class recall is the same as class accuracy..
+On that note, we are able to precisely describe the surface over-prediction and under-prediction issue, by decoupling the problem into which output class that a ground-truth class is most wrongly predicted as, suggesting the model’s inference bias towards an output class given a ground-truth class.
+In the AGNews example, Business is what ground-truth Technology is most wrongly predicted as, suggesting the model’s inference bias to predict Technology as Business.
+To quantify this inference bias, we first conceptualize odd classes.
+An odd class is an output class at inference time which holds the most wrong predictions of a ground-truth class $c_{i}$.
+That is, an odd class is always relative to a ground-truth class.
+For $N$ classes and $M$ test instances, let $y$ and $\hat{y}$ denote the ground-truth class and the prediction respectively, and let $u_{c_{i},c_{j}}$ denote the number of ground-truth $c_{i}$ instances that are wrongly predicted as $c_{j}$,
+$u_{c_{i},c_{j}}=\displaystyle\sum\nolimits_{m}\mathbbm{1}\{y_{m}=c_{i},\hat{y}_{m}=c_{j}\},$ (1)
+$odd_class|(c_{i}):=\operatorname*{arg\,max}_{c_{j}}u_{c_{i},c_{j}},$ (2)
+[…]
+For this goal, it could be direct to use the rate of wrong predictions in an odd class as a measurement of its “hurt” to the accuracy of a ground-truth class, but minimizing these rates alone may not adequately mitigating class accuracy imbalances, which is an ultimate goal of reducing over-/under-predictions (More analyses on this measurement are in Appendix B.).
+A simpler yet better measurement is the absolute accuracy difference between the odd class and the ground-truth class.
+For example, knowing that Business is odd to Tech, the 78% accuracy difference between Business (97%) and Tech (19%) can serve as a measurement of how much Tech is under-predicted.
+The 19% accuracy of Tech also implicitly reflects the many predictions in its odd class Business.
+It should be noted that measuring these accuracy differences sometimes fails to capture the hidden ground-truth-and-odd relationships.
+For instance, World is the odd class to ground-truth Business, but the -12% accuracy difference can not utter this fact.
+Despite this drawback, directly minimizing the accuracy differences will in return significantly reduce predictions in odd classes and improve under-predicted classes’ accuracies.
+The Contextual Oddity Bias.
+To this end, we formulate the definition of Contextual Oddity Bias (COBias).
+First, the absolute value of the accuracy difference between a ground-truth class and its odd class is termed as COBiassingle.
+COBiassingle measures a model’s tendency to over-predict an answer that is caught as odd than the correct answer (or equivalently, a tendency to under-predict the correct answer).
+Notably, for LLMs, an odd answer may be a manifest of surface patterns from prompts or biases in the pretraining data.
+In addition, the odd classes in a same downstream task can vary across different models.
 That’s why we use “contextual” to mean that the oddity of an answer depends on the model or task.
-In addition, the prefix “CO” in COBias can be interpreted as the pairwise nature of the bias, suggesting that if there is an odd class, there has to be a class that is biased to it.
-The Contextual Oddity Bias Metric.
-We define the contextual oddity bias metric as follows.
-Formally, given an $N$-way classification task, let $x_{m}$ denote an input whose label is $y_{m}$.
-By few-shot prompting, let $\boldsymbol{\boldsymbol{p}_{m}}=(p_{m1},\dots,p_{mN})$ denote the output token probability over $N$ classes $c_{1},\dots,c_{N}$.
-The ICL prediction $\hat{y}_{m}$ is $\operatorname*{arg\,max}_{j\in\{1,\dots,N\}}\boldsymbol{p}_{m}$.
-For a given class $c_{i}$, we denote the class it is most biased towards as its oddest class $c_{o}$.
-Suppose test accuracy for the two is $A_{o}$ and $A_{i}$ respectively, then COBiassingle is computed by:
-$\textrm{COBias}_{\textrm{single}}=\displaystyle\sum_{i\neq o}\bigg{|}A_{o}-A_{i}\bigg{|}$ (1)
-Class accuracy $A_{i}$ is given by $\frac{1}{|\mathcal{C}_{i}|}\sum_{m\in\mathcal{C}_{i}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\}$, where $\mathcal{C}_{i}$ is the set of indices of class $c_{i}$ examples, and $\mathbbm{1}(\boldsymbol{\cdot})$ returns 1 if the condition inside is satisfied and 0 otherwise.
-However, bedsides the oddest class $c_{o}$, there could be the second oddest, the third oddest, or etc.
-class for a given class.
-Generalizing the definition, we extend the contextual oddity bias metric as:
-$\textrm{COBias}={N\choose 2}^{-1}\displaystyle\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\bigg{|}A_{i}-A_{j}\bigg{|}$ (2)
-The contextual oddity bias metric reflects the inter-class imbalance in accuracy.
-This metric considers each class as the odd class to others, representing the difference between every pair of classes.
-The Debiasing Problem. The problem is two-fold - we need to reduce the contextual oddity bias, while maintaining or increasing the overall test accuracy.
+Formally, given an $N$ classes and $M$ instances, let $x_{m}$ denote an instance whose label is $y_{m}$.
+By few-shot prompting, the model’s output token probability over $N$ classes $c_{1},\dots,c_{N}$ is $\boldsymbol{p}_{m}=(p_{m1},\dots,p_{mN})$.
+$\boldsymbol{p}_{m}$ is normalized over $N$ classes, $\sum_{n=1}^{N}p_{mn}=1$, $m\in\{1,\dots,M\}$.
+Then the prediction $\hat{y}_{m}$ is $\operatorname*{arg\,max}_{j\in\{1,\dots,N\}}\boldsymbol{p}_{m}$.
+For a given class $c_{i}$, let $A_{\textrm{odd}_{i}}$ and $A_{c_{i}}$ denote the accuracies for the odd class $\textrm{odd}_{i}$ and $c_{i}$ respectively, the COBiassingle metric is computed as:
+$\textrm{odd}_{i}=\verb|odd_class|(c_{i}),$ (3)
+$\textrm{COBias}_{\textrm{single}}=\frac{1}{N}\displaystyle\sum\nolimits_{i}\bigg{|}A_{\textrm{odd}_{i}}-A_{c_{i}}\bigg{|},$ (4)
+where class accuracy $A_{c_{i}}=\frac{1}{|\mathcal{S}_{c_{i}}|}\sum\nolimits_{m\in\mathcal{S}_{c_{i}}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\}$, where $\mathcal{S}_{c_{i}}$ is the set of indices for class $c_{i}$ instances.
+Similarly, $A_{\textrm{odd}_{i}}=\frac{1}{|\mathcal{S}_{\textrm{odd}_{i}}|}\sum\nolimits_{m\in\mathcal{S}_{\textrm{odd}_{i}}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\}$, where $\mathcal{S}_{\textrm{odd}_{i}}$ is the set of indices for class $\textrm{odd}_{i}$ instances.
+Moreover, besides the odd class, which is the most-frequent wrongly predicted class, there could also be a secondly most-, or third most-, to the least-frequent wrongly predicted class.
+For a ground-truth class, there are the odd, the second odd, to the least odd classes.
+For example, for ground-truth Tech (1240 instances), Business is the odd (822), while World is the second odd (156), and Sports is the least odd (27).
+Then, we extend the accuracy difference between the single odd class and a ground-truth class to every pair of classes, and absolute value can be taken.
+This arrives at the ultimate the COBias metric as follows.
+$\textrm{COBias}={N\choose 2}^{-1}\displaystyle\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\bigg{|}A_{c_{i}}-A_{c_{j}}\bigg{|}.$ (5)
+This COBias metric directly reflects the inter-class imbalance in accuracy.
+Indeed, the absolute value of accuracy differences of some pairs in Equation 5 may mask the most frequent wrongly predicted class (the odd class).
+Nonetheless, it reflects a well-rounded measurement of all pairs of classes, whether involving odd classes or not, so it is a powerful metric when we strongly demand minimal class accuracy imbalance.
+In a nutshell, the debiasing problem is minimizing COBias (Eq.
+5) while maintaining or increasing the overall inference accuracy.
+This post-hoc bias evaluation metric will enable easy-to-operate inference-time/post-hoc output likelihood optimization, lifting the heavy computations required for training-wise data selection or model parameter updates.
 3 Improving LLM Outputs by Debiasing as Nonlinear Integer Programming
@@
 The upper-left plot shows that Loc.
-is an odd class to the others, holding their majority wrong predictions, especially for Ent.; the upper-right plot demonstrates the stark contrast in per-class accuracy, with 100% for Loc.
+is an odd class to the others, holding their majority wrong predictions, especially for Ent.; the upper-right plot demonstrates the stark contrast in class accuracy, with 100% for Loc.
 and 0 for Ent..
@@
 In this work, we approach the debiasing problem as a nonlinear integer programming (NIP) problem.
-The objective is to optimize correction weights to rectify the biased, per-class probabilities, with directly modeling the COBias metric and two accuracy-related metrics.
-The weights are optimized on a set of labeled examples, and are simply plugged in to inference-time probabilities.
+The objective is to optimize correction weights to rectify the biased, class probabilities, by directly minimizing the COBias metric while maximizing two accuracy-related metrics.
+The weights are optimized on a set of labeled examples, and are simply plugged in during test time.
 An overview of the proposed DNIP method is shown in Figure 1.
-Our source code will be released.
 3.1 The DNIP Objective
-Considering $M$ examples with known class labels, we let $\boldsymbol{p}_{m}$ denote the normalized output probability over $N$ class tokens for an example, and $\boldsymbol{p}_{m}=(p_{m1},\dots,p_{mn},\dots,p_{mN})$, $\sum_{n=1}^{N}p_{mn}=1$, $m\in\{1,\dots,M\}$.
 To debias $\boldsymbol{p}_{m}$, we introduce a discrete $K$-point correction scale $\boldsymbol{\omega}$ ranging from 0 to 1, to re-weight class probability.
@@
 An example $10$-point scale is given by $\boldsymbol{\omega}=(0.1,\dots,1.0)$.
-For each class probability in $\boldsymbol{p}_{m}$, it times a correction weight chosen from $\boldsymbol{\omega}$ to obtain a debiased class probability.
+For each class probability in $\boldsymbol{p}_{m}$, it multiplies by a correction weight chosen from $\boldsymbol{\omega}$ to obtain a debiased class probability.
 Finding optimal correction weights is a combinatorial optimization problem: for $N$-way class probability $\boldsymbol{p}_{m}$, each class selects a weight from the $K$-dimensional correction scale.
@@
 Instead, we model the debiasing problem as a nonlinear integer program, and search for an optimal set of class correction weights with simulated annealing.
-To this end, the problem boils down to selecting optimal weight indices, and we define an integer selection variable $\xi=(\xi_{1},\dots,\xi_{N})$ that selects a weight index $k$ for the $n$-th class probability in $\boldsymbol{p}_{m}$, namely:
+Without loss of generality, we boil down the problem to selecting optimal weight indices towards minimizing the pairwise class accuracy imbalance, while maximizing the overall accuracy and the pointwise mutual information between class labels and predictions.
+Formally, we define an integer selection variable $\xi=(\xi_{1},\dots,\xi_{N})$ that selects a weight index $k$ for the $n$-th class probability in $\boldsymbol{p}_{m}$, namely:
 $\xi_{n}=k\quad\text{if the $n$-th class selects }\omega_{k}.\$ (6)
@@
 $\xi_{n}\in\{1,\dots,K\},\qquad n=1,\dots,N$ (11)
-where $\beta$ and $\tau$ are parameters that control the bias and the negative PMI terms respectively, $\hat{S}^{j}$ and $S^{j}$ denote the examples of prediction $j$ and true label $j$ respectively [^1] In details, $f(\hat{S}^{j})$ is the ratio between the number of examples with prediction $j$ and the total number of examples, similarly, $f({S}^{j})$ is the ratio between the number of examples with class label $j$ and the total number of examples, $f(\hat{S}^{j},S^{j})$ is the ratio between the number of correct predictions of class $j$ and the total number of examples.
+where $\beta$ and $\tau$ are parameters that control the bias and the negative PMI terms respectively, $\hat{S}^{j}$ and $S^{j}$ denote the examples of prediction $j$ and true label $j$ respectively.
+In details, $f(\hat{S}^{j})$ is the ratio between the number of examples with prediction $j$ and the total number of examples, similarly, $f({S}^{j})$ is the ratio between the number of examples with class label $j$ and the total number of examples, $f(\hat{S}^{j},S^{j})$ is the ratio between the number of correct predictions of class $j$ and the total number of examples.
 Therefore, $\textrm{PMI}_{j}=\log\big{(}\frac{\mathrm{n}(\hat{S}^{j},S^{j})/N}{\mathrm{n}(\hat{S}^{j})/N\cdot\mathrm{n}(S^{j})/N}\big{)}$, where $\mathrm{n}(\cdot)$ is the count.
 In actual computations, we apply add-$u$ smoothing to smooth out zeros in the numerator and denominator.
-The value of the smoothing coefficient $u$ is selected along other parameters on a development set..
+The value of the smoothing coefficient $u$ is selected along other parameters on a development set.
 The debiased class probabilities and prediction are:
@@
 As the second term COBias has minimized the inter-class accuracy bias, the first term error rate brings the class prediction closer to the actual class, and the third term negative PMI further enforces the class prediction to be more confident when it is close to the actual class.
+The PMI term can also be seen as a constraint to penalize each individual class.
 3.2 Solving DNIP with Simulated Annealing
@@
 The computational complexity for the SA algorithm is $\mathcal{O}(NK)$.
-Please refer to the Appendix Section A for derivations.
+Please refer to Appendix C for derivations.
 4 Experiments
@@
 We evaluate three open-source LLMs across varied scales and model families, including GPT-2-XL (1.5B parameters), Llama-2-7B (7B parameters), and Llama-2-13B (13B parameters).
-For ICL, we use 1-shot prompting, and obtain the output per-class softmax probabilities.
+For main results, we use 1-shot prompting and obtain the ICL output class softmax probabilities.
 To standardize calculations, we normalize the softmax probabilities over all classes.
@@
 All prompting is done on an NVIDIA A100 GPU.
-The simulated annealing algorithm executes on CPU (execution time depends on the optimization set size, 3 to 30 minutes on AMD EPYC 7742 CPU or slightly longer on local laptop CPU).
+Simulated annealing executes on CPU (execution time depends on the optimization set size, 3 to 30 minutes on AMD EPYC 7742 CPU or slightly longer on local laptop CPU).
 Model Eval. Metric AGNews DBpedia SST-5 TREC RTE DDI PubMedaQA Avg.
@@
 88.61.7
-16.23.6
+16.23.7
 44.94.3
@@
 14.21.3
-74.20.7
-4.33.4
+74.30.8
+4.33.3
 40.46.0
@@
 The average test accuracy (in black) and COBias (in blue) with standard deviation over three runs are shown in Table 3.
-We find that test COBias by ICL is large in all three LLMs, with GPT-2-XL having the largest average COBias of 50.3%, showing that LLMs of different sizes and families exhibit large per-class accuracy differences.
-With DNIP, the average test COBias over three models reduces from 43% to 16%, and the test accuracy increases from 52% to 64%.
-Specifically, DNIP achieves an 16%, 12%, and 10% absolute improvement in test accuracy for GPT-2-XL, Llama-2-7B, and Llama-2-13B respectively; meanwhile, DNIP significantly reduces test COBias by an absolute 29%, 23%, and 26% for the three models, over ICL results.
-As for detailed results, for GPT-2-XL where the ICL test accuracy is below 30% (DDI), DNIP boosts test accuracy by 29% while lowering test COBias from 41% to 17%; for Llama-2 7B and 13B where the ICL test accuracy is already in higher 80s, e.g., DBpedia, DNIP can further improve the accuracy to 90s while reducing COBias.
-In addition, we observe a very large ICL test COBias at 82.4% for GPT-2-XL on RTE.
+In all three LLMs, test COBias by ICL is large, with GPT-2-XL exhibiting the largest COBias (50.3%), showing that class accuracy imbalance exists in LLMs of different sizes and families.
+With DNIP, test COBias is significantly reduced, by an absolute 29%, 23%, and 26% for the three models respectively; meanwhile, there is an 16%, 12%, and 10% absolute improvement in test accuracy respectively.
+It is worth noting that for tasks whose ICL test accuracy is already in higher 80s, e.g., DBpedia, DNIP can reduce COBias while further improving the accuracy to 90s.
+In addition, we observe a very large ICL COBias at 82.4% in GPT-2-XL for RTE, reflecting severe over-predictions in class True.
 After applying DNIP, the COBias greatly reduces to 7.1%.
-The 82.4% COBias on RTE suggests that the conventional ICL predictions mostly fall in a single class, failing to predict the other class.
-In fact, The RTE task seems too challenging that GPT-2-XL’s output probability of class True is much higher than that of class False.
-Though re-weighting can lower both probabilities, it may not gain a corrected prediction.
-Still, the over 75% absolute reduction in COBias demonstrates that DNIP in the worst case can balance severely skewed ICL predictions without hurting the overall accuracy much.
-On two biomedical tasks, DNIP brings an average absolute increase of 34% and 11% in accuracy, and an average absolute reduction of 31% and 13% in COBias, showing its effectiveness for domain-specific tasks.
+That is, DNIP obtains over 75% absolute reduction in COBias for RTE, demonstrating DNIP’s capability for balancing very imbalanced class accuracies without hurting the overall accuracy much.
+Moreover, on the two biomedical tasks, DNIP brings an average absolute reduction of 31% and 13% in COBias, and an average absolute increase of 34% and 11% in accuracy, showing its effectiveness for domain-specific tasks.
+Moreover, the 13B Llama model can suffer even higher COBias than its 7B counterpart on biomedical tasks, further showing that class accuracy imbalance does not go away as model sizes increase, and DNIP is effective to balance class accuracies for LLMs of all sizes.
+Back to the AGNews example in Table 2, we present the DNIP debiased results in Table 2.
+Over-predictions in Business are greatly reduced, and the accuracy of Tech is significantly improved from 19% to 84%.
 These results show that DNIP strongly mitigates COBias and greatly improves the accuracy of lower-performing classes, while boosting overall accuracy.
-Moreover, relatively large LLMs can suffer even higher COBias than smaller models on biomedical tasks, suggesting that the per-class accuracy bias does not go away as models scale, and debiasing is essentially needed to improve inference for larger LLMs.
-4.3 DNIP Balances Class Accuracies across Varied Classification Tasks, as many as 14 Classes
-Figure 2: Comparisons of pairwise class accuracy differences on test sets before and after applying DNIP; the pinker the higher difference, the greener the lower difference; average accuracy over three runs are used.
-Heatmaps after debiasing show nearly 0 differences for many class pairs, demonstrating DNIP’s effectiveness for COBias reduction.
-We visualize pairwise class accuracy differences before and after applying DNIP on Llama-2-13B by the heatmaps in Figure 2, which show that DNIP is straightforwardly effective in reducing COBias.
-In each heatmap, the value of row $i$, column $j$ cell represents the absolute difference between the test accuracies of two classes $i,j$ , i.e., $|A_{i}-A_{j}|$.
-We observe that DNIP brings the accuracy differences down to close to zero for 4 out of 7 datasets, including AGNews, DDI, DBpedia, and RTE; and the most relative reduction is seen on TREC.
-These results show that DNIP consistently reduces prediction bias across different NLP classification tasks.
-Figure 3: Test confusion matrix for DBpedia before and after applying DNIP. DNIP greatly reduces wrong predictions for Nat., which is the lowest-accuracy class before debiasing.
+Computational Time.
+Regarding computational time needed to perform DNIP optimization, on all datasets, the optimization is finished in the scale of minutes.
+Though computational time increases with more classes and a more fine-grained weight scale, it should be noted that the time elapsed in optimization is dozens of minutes even for tasks of more than 10 classes, and a weight scale of 30 is good enough for most of the datasets experimented in this paper.
+This demonstrates the practicality of DNIP in the computational time needed for performing optimization.
+Prompting Method Metric AGNews DBpedia SST-5 TREC RTE DDI PubMedQA Avg.
+Acc $82.5_{2.0}$ $93.6_{1.3}$ $45.8_{6.0}$ $58.7_{23.3}$ $61.9_{16.9}$ $34.1_{42.1}$ $70.4_{7.1}$ 63.9
+5-shot ICL COBias $24.4_{4.2}$ $9.0_{2.0}$ $48.0_{15.4}$ $35.4_{13.7}$ $61.3_{40.4}$ $44.4_{5.4}$ $52.2_{19.3}$ 39.2
+Acc $88.5_{0.6}$ $95.8_{0.7}$ $52.9_{2.5}$ $76.6_{8.7}$ $75.8_{4.5}$ $54.1_{5.1}$ $59.9_{1.4}$ 71.9
+5-shot DNIP COBias $7.0_{0.9}$ $5.7_{1.1}$ $15.8_{12.6}$ $14.4_{7.5}$ $3.1_{1.9}$ $16.5_{7.6}$ $9.6_{4.6}$ 10.3
+Acc $83.5_{1.5}$ $95.2_{1.2}$ $50.3_{2.3}$ $67.0_{12.7}$ $75.0_{0.8}$ $9.7_{1.0}$ $52.3_{5.3}$ 61.9
+k-shot (1 demon. per class) ICL COBias $14.9_{5.1}$ $7.0_{2.2}$ $36.3_{7.2}$ $38.2_{5.1}$ $22.5_{13.2}$ $39.7_{3.5}$ $20.9_{4.2}$ 25.6
+Acc $88.7_{0.5}$ $96.6_{0.5}$ $51.3_{1.0}$ $82.7_{1.4}$ $76.7_{3.7}$ $43.6_{6.1}$ $55.3_{2.6}$ 70.7
+k-shot (1 demon. per class) DNIP COBias $7.3_{0.5}$ $4.3_{0.7}$ $2.8_{1.5}$ $12.1_{5.5}$ $5.0_{3.3}$ $12.5_{4.3}$ $8.7_{1.2}$ 7.5
+Table 4: DNIP results in more ICL settings. Average score over three runs are reported.
+4.3 More ICL Settings
+We prompted Llama-2-13B in two additional settings: 5-shot, where each demonstration is randomly selected as in the 1-shot case; k-shot, where k is the number of classes and each class is represented by a demonstration example in the prompt.
+Table 4 presents results on seven benchmark datasets in both settings.
+For example, the relative average COBias reduction (comparing DNIP to ICL) is 65% and 74% for 1-shot and 5-shot cases respectively, and the relative average accuracy increase is 16% and 11% for 1-shot and 5-shot cases respectively, demonstrating DNIP’s capabilities with more shots.
+In addition, prompt design does contribute to different output probabilities, which could be helpful to provide different starting points for DNIP, so DNIP could optimize for better solutions.
+However, prompt engineering alone may not be most effective to solve the COBias issue.
+For example, compared to 1-shot prompting (Table 3), increasing the number of shots does not significantly further reduce COBias or boost accuracy; using more diverse demonstrations (k-shot) may help with COBias, but does not gain much higher accuracy than 1-shot.
+Therefore, using more shots does not help reduce COBias much, and DNIP significantly reduces COBias and improves accuracy in both 5-shot and k-shot settings, further showcasing its effectiveness.
+4.4 DNIP Significantly Reduces Over-predictions, and Balances Class Accuracies across Varied Classification Tasks, as many as 14 Classes
+Figure 2: Pairwise class accuracy differences on test sets before and after applying DNIP; the pinker the higher class accuracy difference, the greener the lower; average accuracy over three runs are used.
+Heatmaps after DNIP show nearly 0 accuracy differences for many class pairs.
+Figure 3: Test confusion matrix for DBpedia before and after applying DNIP. DNIP greatly improves ICL predictions for Nat., which is the most under-predicted class by ICL.
+We visualize pairwise class accuracy differences before and after applying DNIP to 1-shot, Llama-2-13B ICL results in Figure 2.
+In each heatmap, row $i$, column $j$ cell represents the absolute difference between the test accuracies of two classes in row $i,j$ , i.e., $|A_{i}-A_{j}|$.
+DNIP brings the accuracy differences down to close to zero for AGNews, DDI, DBpedia, and RTE.
+The lighter green heatmaps in the bottom row show that DNIP consistently reduces COBias across different NLP classification tasks.
 Relieving Odd Classes for 14-Way Ontology Classification.
@@
 Before debiasing, the Nat.
-class exhibits the largest class accuracy difference compared to other classes, suggesting that Nat.
-is either the highest-accuracy class manifesting a strong prediction bias towards Nat., or the lowest-accuracy class where predictions of true Nat.
-examples are biased to one or more odd classes.
-We plot the confusion matrix of test results of 1 run out of 3 runs in Figure 3 and find that Nat.
-has the lowest accuracy, and its predictions are biased to two odd classes, Trans.
-and Bldg., with Trans.
-being the oddest to Nat..
+class exhibits the largest class accuracy difference to other classes.
+In fact, Nat.
+is the most under-predicted class, whose odd class is Trans.
+and second odd class is Bldg., as shown by the confusion matrix in Figure 3 (based on seed 0, 1-shot ICL run).
 DNIP greatly reduces false negatives for Nat., i.e., reduces wrong Trans.
@@
 predictions for Nat..
-Besides these visualizations, on an average over three runs, the test accuracy for Nat.
+Moreover, on an average over all three runs, the test accuracy for Nat.
 improves from 44% to 89%, and that for Trans.
-also increases from 78% to 87%, demonstrating DNIP’s effectiveness in reducing pairwise class accuracy bias for classification tasks of as many as 14 classes.
-Furthermore, we see more balanced class accuracies with DNIP for smaller LLMs of Llama-2-7B and GPT-2-XL, in Appendix Section B.
-4.4 DNIP Objective Ablation
-We present ablation analysis on objective functions to show that the integration of accuracy-related objectives and the COBias objective are indispensable in achieving the best of both worlds.
-Recall three parts in our objective function $z$: the error rate term $z_{1}=\frac{1}{M}\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}$; the COBias term $z_{2}={N\choose 2}^{-1}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}|A_{i}-A_{j}|$; and the PMI term $z_{3}=\sum_{j=1}^{N}\textrm{PMI}_{j}$; the accuracy-related terms are $z_{1}$ and $z_{3}$.
-We set up 7 different objective functions to perform DNIP: $z_{1},z_{2},z_{3},z_{1}+\beta z_{2},z_{1}-\tau z_{3},\beta z_{2}-\tau z_{3},z_{1}+\beta z_{2}-\tau z_{3}$, and compare their effectiveness on Llama-2-13B in Figure 4.
-Ablations on GPT-2-XL and Llama-2-7B show similar findings (Appendix Section C).
+also increases from 78% to 87%, demonstrating that DNIP is effective in reducing class accuracy differences for tasks of as many as 14 classes.
+Furthermore, we present more heatmaps for smaller LLMs including Llama-2-7B and GPT-2-XL in Appendix D.
+4.5 Optimizing COBiassingle
+Using COBiassingle (Eq. 4) for DNIP optimization will focus on odd classes only. Figure 4 shows ablation results when replacing the COBias term in the DNIP objective with COBiassingle.
+$\min\quad Z(\xi)=\frac{1}{M}\displaystyle\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}+\beta\frac{1}{N}\displaystyle\sum\nolimits_{i}\bigg{|}A_{\textrm{odd}_{i}}-A_{c_{i}}\bigg{|}-\tau\displaystyle\sum_{j=1}^{N}\textrm{PMI}_{j}$ (14)
+Figure 4: Test COBiassingle, COBias, and overall accuracy scores (from left to right).
+Blue dots represent the respective evaluation metric score of the original 1-shot ICL Llama-2-13B outputs of the seven benchmark datasets; green dots are debiased results optimized by the COBiassingle DNIP objective (Eq.
+14); red dots are debiased results optimized by the COBias DNIP objective.
+Comparing DNIP with COBiassingle and DNIP with COBias, their performances are close, but DNIP with COBias achieves higher overall test accuracy and slightly better bias mitigation performances.
+In other words, DNIP with COBiassingle can significantly minimize accuracy differences between odd classes and ground-truth classes, but DNIP with COBias is even more effective when it comes to balancing all pairs of classes, not limited to the odd classes.
+In summary, optimizing either COBiassingle or COBias is an effective inference-time measure to reduce over-predictions in raw ICL outputs and boost accuracy for under-predicted classes.
+4.6 Accuracy-related Objectives vs. the COBias Objective
+We show that the integration of accuracy-related objectives and the COBias objective are indispensable in achieving the best of both worlds in Figure 5.
+DNIP objective function $z$ consists of three terms: the error rate term $z_{1}=\frac{1}{M}\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}$; the COBias term $z_{2}={N\choose 2}^{-1}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}|A_{i}-A_{j}|$; and the PMI term $z_{3}=\sum_{j=1}^{N}\textrm{PMI}_{j}$; the accuracy-related terms are $z_{1}$ and $z_{3}$.
+We ablate all ablation objective functions to perform DNIP: $z_{1},z_{2},z_{3},z_{1}+\beta z_{2},z_{1}-\tau z_{3},\beta z_{2}-\tau z_{3},z_{1}+\beta z_{2}-\tau z_{3}$.
+Ablations on GPT-2-XL and Llama-2-7B show similar findings (Appendix E).
 To make fair comparisons, we fix $\beta=2.7,\tau=0.2$ and weight scale $K=30$, which may not result in the same scores reported in Table 3.
-Figure 4: Ablation analysis of the objective function, exemplified with Llama-2-13B; test scores averaged over 7 datasets are used for the rightmost plot.
 Among all ablations, $z_{1}+\beta z_{2}-\tau z_{3}$ achieves a balance point between accuracy and COBias.
@@
 (4) Using $z_{1}+\beta z_{2}$ or $\beta z_{2}-\tau z_{3}$ is promising with a similar/lower accuracy and a slightly lower COBias.
-We note that accuracy-related objectives alone increase accuracy by over 16% ($75\%-59\%$), which is greater than using the objective combining COBias, suggesting the NIP can be a direction for those who look for accuracy improvement.
+We note that accuracy-related objectives alone increase accuracy by over 16% ($75\%-59\%$), which is greater than using the objective combining COBias, suggesting DNIP is effective for those who emphasize accuracy improvement.
+Figure 5: Ablation analysis of the objective function, exemplified with Llama-2-13B; test scores averaged over 7 datasets are used for the rightmost plot.
 4.7 Performance Improvements vs. Number of Optimization Examples
@@
 We further notice a surge in reductions from 1,000 to full for DDI, SST-5, and DBpedia (full optimization sizes: 9,500, 8,116, 9,500), suggesting that around 8,000 examples are needed for a further emergent reduction - PubMedQA may see further COBias reduction if more optimization examples are available.
-Figure 6: Relative test accuracy increase; higher is better.
 Test accuracy is improved with only 10 optimization examples, and plateaus at around 500 examples.
@@
 At 1,000 examples, test accuracy for DDI rises to 52%, which is more than 6 times of the ICL accuracy, showing that DNIP is effective for this challenging biomedical NLP task.
+Figure 7: Relative test accuracy increase; higher is better.
 Moreover, we observe for RTE a declining trend of accuracy increase, which matches its growing trend of COBias reduction, reflecting the trade-off between the two objectives.
@@
 Overall, DNIP improves both accuracy and COBias at 10 examples, and achieves acceptable COBias reduction and accuracy improvement at 100 examples, with test COBias reduced by 57% and test accuracy improved.
-For detailed scores, please refer to Appendix D.
+For detailed scores, please refer to Appendix F.
 5 Related Work
@@
 [23] shows that LLM reasoning abilities are affected by adding biasing features to the input.
-Along this line, we tackle a common LLM failure of the per-class accuracy difference, via nonlinear integer programming optimization on the evaluation metric scores.
+Along this line, we tackle a common LLM failure of the class accuracy difference, via nonlinear integer programming optimization on the evaluation metric scores.
+The inference-time class accuracy bias has multi-faceted causes.
+It could also be attributed to pre-training data, such as predilections for common tokens in the pretraining corpora [24].
+It could be a direct outcome of LLMs’ liking for specific patterns or answers, leading to highly imbalanced predicted class distributions [25].
+Other causes include choices of prompt templates [26, 27, 28], demonstrations [29], and prompt orders [22, 23].
+Unintended LLM behaviors may also exacerbate the bias, such as user-following sycophancy [30, 31, 32] and prompt toning effects [21, 33, 34, 35].
 6 Discussion
+More Methods That Potentially Improve Accuracy while Reducing COBias.
+Prior methods tackle the prediction bias problem from different angles.
+Although COBias is not be explicitly modeled, they are promising to reduce COBias.
+Therefore, we draw comparisons to two popular categories of debiasing methods, adaptations and calibrations.
+In particular, with the same class probability vectors as input, we build a single linear adapter (LA) network with adapter dimension of 64, and train it towards the same objective as DNIP; we also follow the batch calibration method [36] to estimate a calibration term using test examples.
+Figure 8: Comparisons to more methods.
+Figure 8 shows the comparison results. fine-tuning a simple linear adapter does not work well on the class output probabilities. The proposed method outperforms both methods in accuracy and COBias.
 Using Chain-of-Thought Prompting to Improve LLM Predictions.
@@
 Another hazard is using prompts to compute the offset, which may cascade biases brought by a certain or several prompt templates, and biases inherited from pretraining may not be reflected.
-Another line of works does not rely on further prompting.
-Han et al.
-[3] fits prototypical clusters on class probabilities and updates predictions by a cluster-label mapping.
-We share the same spirit of correcting biased predictions without more prompting, but we differ in that our correction weights are explicitly optimized on COBias and accuracy metrics, while the cluster-based decision boundary is implicitly learned.
+Instead, we correcct biased predictions without more prompting, but we differ in that our correction weights are explicitly optimized on COBias and accuracy metrics, while the cluster-based decision boundary is implicitly learned.
 Evaluating Closed LLMs.
@@
 For example, ChatGPT returns log probabilities of up to 20 most likely output tokens, and may not cover the probability for every class that is needed in our class prediction debiasing.
+Optimization Algorithms for Integer Programming.
+The classic solvers for integer programming include operational research techniques such as Branch and Bound (BnB), commonly used for linear integer programming problems.
+However, it is difficult for these methods to solve our DNIP model, which contains discontinuous functions and is non-differentiable.
+In this case, a series of metaheuristic algorithms are more suitable, solversincluding Simulated Annealing (SA), Genetic Algorithms, Ant Colony Optimization, and Particle Swarm Optimization.
+These algorithms all belong to the same category, and each meta heuristic has their own pros and cons.
+Therefore, SA in this paper has a clear purpose of tackling our NIP mathematical model that classical operational research methods cannot solve.
+Our Motivation Is Different from Classical Post-hoc Corrections.
+Some may contend that achieving equitable accuracies across all classes is a well-explored issue in standard machine learning classifiers.
+However, it is crucial to see that class prediction accuracy imbalances must be understood within their specific context.
+The accuracy bias in LLMs’ outputs arises from fundamentally different causes than the unequal class accuracies seen in potentially overfitted traditional classifiers.
+In the former, biases stem from prompts and pre-training, while in the latter, the imbalance is driven by skewness in the supervised training data.
+This distinction underscores why our approach is uniquely applied to LLMs’ output token class probabilities.
 7 Conclusion
-This work tackles the imbalance in per-class prediction accuracy of language models.
-We introduce COBias to assess pairwise class accuracy differences, and reveal that LLMs of varied scales and families exhibit large COBias at inference.
-To minimize COBias and push for higher accuracy, we view debiasing as a combinatorial optimization problem and propose DNIP, which optimizes class correction weights by jointly minimizing evaluation scores by COBias and maximizing accuracy.
+This work tackles the imbalance in class prediction accuracy of language models.
+We introduce COBias and reveal that LLMs of varied scales and families exhibit large COBias at inference.
+To minimize COBias and push for higher accuracy, we view debiasing as an inference-time combinatorial optimization problem and propose DNIP, which optimizes class correction weights by jointly minimizing evaluation scores by COBias and maximizing accuracy.
 Empirical evaluations on three LLMs across seven multi-class NLP classification tasks demonstrate that DNIP simultaneously achieves lower test COBias and higher test accuracy over conventional ICL.
 We advocate that reducing COBias is a direction to gain more accurate, reliable LLM predictions.
-Limitations and Future Work.
-The limitation is that we only evaluate NLP classification tasks.
-We leave evaluations on open-ended generations and other modalities for future work.
-Indeed, DNIP is fruitful for debiasing imbalanced class predictions, which can easily extend to vision/multimodal LMs, as long as there is a need for mitigating per-class accuracy differences.
-This simple inference-time plug-in solution is there for further boosting LM performances.
-Appendix A Derivation of SA Algorithm Complexity
+For future works, we may extend our studies to more tasks and other modalities.
+We aim to find simple inference-time plug-in solution for further boosting language model prediction performances.
+Appendix A Other Examples of Over-prediction and Under-prediction in LLMs
+We present Llama-2-13B’s and ChatGPT’s over- and under-predictions on a class imbalanced classification task, the Drug-Drug Interaction (DDI) relation classification task, where around 90% test instances are Negative.
+As shown in Table 5, Llama-2-13B significantly under-predicts Negative.
+In Table 6, similar under-predictions in the DDI Negative class happens for ChatGPT (GPT-3.5-turbo-instruct [39]), the RLHF’d chat LLM with magnitudes of more parameters.
+(It is noted that at the time of our experiments, ChatGPT does not support returning full probabilities over the entire vocabulary.
+To investigate the over-/under-prediction issue, we prompt ChatGPT on the DDI task with the following prompt template: “Please choose one most suitable answer for the following description.
+[ONE-SHOT DEMONSTRATION] Description: [TEST INPUT] Q: What is the drug-drug interaction relation between the aforementioned @DRUG\$ pair?
+Choose from Negative, Mechanism, Effect, Advice, Interaction.
+A: ”, and use the answer token directly as the prediction, and plot the confusion matrix in Table 6.
+For fair comparisons, we prompt Llama-2-13B in the same manner to obtain Table 5, which is different from how we prompt Llama-2-13B in the main paper.)
+Pred.
+Neg Eff Mech Adv Int
+Neg 1749 438 1498 546 506
+Eff 118 33 110 52 47
+True Mech 77 28 123 45 29
+Adv 84 15 62 36 24
+Int 34 11 28 13 10
+Table 5: Over-prediction and under-prediction in Llama-2-13B, as shown by the confusion matrix on DDI test set.
+Pred.
+Neg Eff Mech Adv Int
+Neg 809 707 1001 217 2003
+Eff 45 109 79 6 121
+True Mech 15 46 172 0 69
+Adv 24 12 18 79 88
+Int 9 9 7 1 70
+Table 6: Over-prediction and under-prediction in ChatGPT, as shown by the confusion matrix on DDI test set.
+The over-/under-predictions in LLMs of different sizes and families, and the inconveniences in training/fine-tuning LLMs, together call for a more actionable bias metric quantifying the over-/under-prediction issue and an accompanying debiasing method tailored for LLMs.
+Appendix B Directly Optimizing the Odd Prediction Rate
+We discuss another option of measuring class accuracy imbalances, which does not mitigate odd predictions as effective as using COBias, though using it can obtain similar or slightly higher overall accuracy.
+It is the odd prediction rate, i.e., the rate of odd predictions over the number of ground-truth instances, which is a measurement reflecting the odd class and the level of “accuracy hurt” caused by the odd class.
+We define $H(c_{i})$ as below.
+$\textrm{odd}_{i}=\verb|odd_class|(c_{i}),$ (15)
+$H(c_{i})=\frac{\displaystyle\sum\nolimits_{m}\mathbbm{1}\{y_{m}=c_{i},\hat{y}_{m}=\textrm{odd}_{i}\}}{\displaystyle\sum\nolimits_{k}\mathbbm{1}\{y_{k}=c_{i}\}},$ (16)
+where $\textrm{odd}_{i}$ is the odd class to $c_{i}$.
+For example, one may quantify the effect of a current odd class Business to the ground-truth class Tech as the rate of Business predictions over ground-truth Tech instances, i.e., 822/1240, and an immediate debiasing goal is to minimize this rate.
+The DNIP objective with the odd prediction rate is as follows.
+$\min\quad Z(\xi)=\frac{1}{M}\displaystyle\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}+\beta\frac{1}{N}\displaystyle\sum\nolimits_{i}H(c_{i})-\tau\displaystyle\sum_{j=1}^{N}\textrm{PMI}_{j}$ (17)
+We compare the effectiveness of using COBias, COBiassingle, and $H$ in the DNIP objective, and refer to the respective objectives as ZCB (Eq.
+4), ZCBs (Eq.
+14), and ZH (Eq.
+17).
+Detailed test scores for the benchmark datasets on Llama-2-13B, including class accuracies, are shown in Tables 7 to 13.
+In short, ZH still leaves some classes under-predicted, as examplified by the individual test class accuracies, and test COBias (or test COBiassingle) scores on TREC, RTE, etc.
+The rationale is that optimizing bias measurements that extend accuracy differences, such as the odd prediction rates, does not mitigate odd class predictions better than optimizing COBias directly.
+ZCB most effectively mitigates the over- and under-prediction issue, balancing class accuracies.
+Optimizing other extended measurements, such as precision or $F_{1}$ scores differences, may obtain similar performances to the three existing metrics, which will be left for future discussions.
+DNIP Objective (Results without DNIP are shown on the last row.) Evaluation Metric
+CB CBs H Overall Acc. Class Acc.
+Co. Sch. Art. Ath. Polit. Trans. Bldg. Nat. Vill. Anim. Plt. Alb. Film Bk.
+ZCB 7.7 7.5 3.3 93.4 82.6 97.7 94.0 99.5 97.9 94.4 86.6 89.0 99.5 96.3 98.5 98.6 97.0 77.8
+ZCBs 9.1 5.7 3.9 92.6 87.4 96.8 97.1 98.9 98.1 87.3 87.6 84.0 99.3 99.2 97.9 99.6 96.5 69.0
+ZH 8.9 9.5 2.8 93.2 76.2 99.3 95.0 99.4 98.8 96.9 82.8 89.8 99.5 98.4 98.9 99.8 96.1 76.3
+ICL 16.2 13.7 5.6 88.6 91.7 99.1 92.7 99.7 97.0 94.8 78.4 44.2 99.7 96.9 99.3 99.8 95.2 63.3
+Table 7: Comparison of DBpedia test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
+DNIP Objective (Results without DNIP are shown on the last row.) Evaluation Metric
+CB CBs H Overall Acc. Class Acc.
+Wrd. Sports Bus. Tech.
+ZCB 6.3 4.4 7.2 87.9 85.3 95.9 86.5 83.9
+ZCBs 7.3 4.3 6.9 88.4 84.8 98.2 85.8 84.7
+ZH 8.5 4.8 6.5 88.3 86.8 98.7 84.2 83.2
+ICL 28.3 22.5 14.6 79.9 81.5 98.7 91.0 48.0
+Table 8: Comparison of AGNews test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
+DNIP Objective (Results without DNIP are shown in the last row.) Evaluation Metric
+CB CBs H Overall Acc. Class Acc.
+Neg. Eff. Mech. Adv. Int.
+ZCB 7.5 7.0 39.6 40.4 42.1 31.6 32.8 33.9 31.6
+ZCBs 19.9 16.8 42.1 22.9 19.9 34.2 37.3 39.8 44.8
+ZH 43.6 44.6 30.9 21.7 17.6 12.6 47.2 85.5 33.3
+ICL 45.6 71.8 59.4 7.2 0.0 30.7 61.5 24.1 43.5
+Table 9: Comparison of DDI test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
+DNIP Objective (Results without DNIP are shown in the last row.) Evaluation Metric
+CB CBs H Overall Acc. Class Acc.
+Num. Loc. Per. Desc. Ent. Ab.
+ZCB 14.2 16.0 18.9 77.1 80.2 92.2 88.7 69.1 64.5 74.1
+ZCBs 15.0 13.6 17.7 79.0 90.0 89.3 87.7 70.0 64.5 74.1
+ZH 21.3 31.7 12.6 77.3 93.8 88.5 91.3 75.6 39.7 88.9
+ICL 35.9 19.9 18.5 68.5 82.9 89.7 88.7 72.9 9.9 96.3
+Table 10: Comparison of TREC test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
+DNIP Objective (Results without DNIP are shown in the last row.) Evaluation Metric
+CB CBs H Overall Acc. Class Acc.
+False True
+ZCB 4.5 4.5 26.0 74.1 71.8 76.3
+ZCBs 4.5 4.5 26.0 74.1 71.8 76.3
+ZH 13.3 13.3 24.7 75.7 68.7 82.0
+ICL 43.4 43.4 29.7 71.5 48.6 92.0
+Table 11: Comparison of RTE test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
+DNIP Objective (Results without DNIP are shown in the last row.) Evaluation Metric
+CB CBs H Overall Acc. Class Acc.
+Trbl. Bad Ok Gd. Grt.
+ZCB 18.7 13.2 36.8 48.3 47.0 56.6 19.4 54.1 56.6
+ZCBs 29.3 22.5 38.6 49.3 50.7 59.3 0.9 68.9 54.4
+ZH 35.6 34.8 34.8 50.6 37.2 70.9 0.9 68.8 53.3
+ICL 53.1 63.9 53.1 44.9 28.7 71.6 0.0 81.6 10.8
+Table 12: Comparison of SST-5 test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
+DNIP Objective (Results without DNIP are shown in the last row.) Evaluation Metric
+CB CBs H Overall Acc. Class Acc.
+Yes No Mayb.
+ZCB 41.1 31.3 38.9 63.1 74.4 61.1 12.7
+ZCBs 16.1 17.1 37.1 52.4 56.9 51.5 32.7
+ZH 36.3 24.3 27.8 62.7 65.9 72.0 17.6
+ICL 61.2 65.0 44.0 55.1 42.2 93.7 1.8
+Table 13: Comparison of PubMedQA test set evaluation scores (%) for three DNIP objective functions ZCB, ZCBs, and ZH. Average scores over three runs of different 1-shot demonstrations are reported.
+Appendix C Derivation of SA Algorithm Complexity
 […]
@@
 […]
-Appendix B Additional Heatmaps for Pairwise Accuracy Differences
-Figure 7 and 8 demonstrate that more balanced class accuracies are achieved with DNIP for smaller LLMs of Llama-2-7B and GPT-2-XL.
+Appendix D Additional Heatmaps for Pairwise Accuracy Differences
+Figure 9 and 10 demonstrate that more balanced class accuracies are achieved with DNIP for smaller LLMs of Llama-2-7B and GPT-2-XL.
 Figure 9: Comparisons of pairwise test class accuracy differences before and after applying DNIP on GPT-2-XL; the pinker the higher accuracy difference, the greener the lower accuracy difference; average accuracy over three runs are used.
 Figure 10: Comparisons of pairwise test class accuracy differences before and after applying DNIP on Llama-2-7B; the pinker the higher accuracy difference, the greener the lower accuracy difference; average accuracy over three runs are used.
-Appendix C Additional Ablation Analysis
-We show ablation analysis on GPT-2-XL and Llama-2-7B in Figure 9 and 10.
+Appendix E Additional Ablation Analysis
+We show ablation analysis on GPT-2-XL and Llama-2-7B in Figure 11 and 12.
 Figure 11: Ablations on objective function, GPT-2-XL, demonstrating that $z_{1}+\beta z_{2}-\tau z_{3}$ achieves a balance point between accuracy and COBias.
 Figure 12: Ablations on objective function, Llama-2-7B, demonstrating that $z_{1}+\beta z_{2}-\tau z_{3}$ achieves a balance point between accuracy and COBias.
-Appendix D Full Table with Varying Number of Optimization Examples for DNIP
-The full results for all three models across varying optimization set sizes are shown in Table 2 (on the next page).
+Appendix F Full Table with Varying Number of Optimization Examples for DNIP
+The full results for all three models across varying optimization set sizes are shown in Table 14.
 Opt. set Eval. Metric AGNews DBpedia SST-5 TREC RTE DDI PubMedaQA
@@
 88.61.7
-16.23.6
+16.23.7
 44.94.3
@@
 14.21.3
-74.20.7
-4.33.4
+74.30.8
+4.33.3
 40.46.0

### v5

@@
-Balancing Class Accuracies for Language Models in Inference Time via Nonlinear Integer Programming
+COBias and Debias: Balancing Class Accuracies for Language Models in Inference Time via Nonlinear Integer Programming
 Abstract
-Large language models are good knowledge bases but struggle to perform equally well for all classes in simple text classification tasks.
+Large language models (LLMs) are good knowledge bases but struggle to perform equally well for all classes in text classification tasks.
 This paper investigates a fundamental inference-time problem in language models: imbalanced class accuracies.
-We find what’s underneath the issue is a tendency to over-predict some classes while under-predicting some others, for both base LLMs and more sophisticatedly trained, larger chat LLMs.
-This class accuracy imbalance is difficult to solve from the root via “better” pre-training or fine-tuning strategies, but we show it can be effectively mitigated via inference-time optimization.
+We find what’s underneath the issue is a tendency to over-predict some classes while under-predicting some others.
+This class accuracy imbalance is difficult to solve from the root via better pre-training or fine-tuning strategies, but we show it can be effectively mitigated via inference-time combinatorial optimization.
 To this end, we conceptualize and quantify the over- and under-prediction issue as the Contextual Oddity Bias (COBias), and propose the Debiasing as Nonlinear Integer Programming (DNIP) model to correct in-context learned class probabilities based on minimizing COBias and maximizing overall accuracy, without LLM parameter update.
-Considering that the DNIP model implicitly contains non-differentiable elements, we therefore use the simulated annealing algorithm to solve.
-We perform extensive evaluations on three LLMs across seven NLP classification tasks, in different prompting settings.
-Results show that DNIP simultaneously achieves significant COBias reduction (-27%) and accuracy improvement (+12%) over the conventional ICL approach, suggesting that inference-time mitigation of class accuracy imbalance is a promising direction to push forward LLM performances.
+Considering that the DNIP model implicitly contains non-differentiable elements, we therefore use the simulated annealing algorithm to solve it.
+Extensive evaluations on three LLMs across seven NLP classification tasks in different prompting settings show that DNIP simultaneously achieves significant COBias reduction (-27%) and accuracy improvement (+12%) over the conventional ICL approach, suggesting that inference-time mitigation of class accuracy imbalance is a promising direction to push forward LLM performances.
 1 Introduction
 The fact that large language models (LLMs) achieve remarkable overall test accuracy by few-shot in-context learning (ICL) [1] can obscure a large difference in individual class accuracies.
-Mitigating class accuracy imbalance is under-explored, but important to facilitate a smoother user experience.
-Not mitigating such imbalances may lead to serious damages in some cases.
-For example, a patient can trust an LLM about drug-drug interaction type predictions without knowing that the model over-predicts “no interaction” and has low accuracies for other categories of interactions.
-If the patient trusts “no interaction” when the ground truth is an actual drug interaction type, taking the drugs together can put the patient’s life at risk.
-Therefore, it is high time we took actions to reduce class accuracy imbalances, better without hurting overall accuracy.
-It is difficult to disentangle all the factors that cause class accuracy imbalance, including biases incurred in pre-training, supervised fine-tuning, RLHF, prompting specifics, etc.
-Therefore, it is challenging to comprehensively resolve the root causes through pre-training or fine-tuning methods, or approaches like data selection or data augmentation.
-To overcome these difficulties, we conceptualize a post-hoc bias metric, the Contextual Oddity Bias (COBias; Section 2), to quantify the class accuracy imbalance based on inference-time over-predictions and under-predictions.
-To reduce COBias and rectify LLM outputs, we directly operate on the class probabilities obtained from in-context learning (ICL).
-We propose DNIP, the Debiasing as Nonlinear Integer Programming model, to adjust class probabilities towards lower COBias and higher overall accuracy, from a combinatorial optimization viewpoint (Section 3).
-We formulate the debiasing problem as a discrete correction weight selection problem.
-In particular, DNIP re-weights class probability with an optimal set of weights selected based on both metrics of accuracy and COBias, and thus obtains much less biased predictions with higher accuracy.
-On three widely used different-scale LLMs across seven NLP evaluation datasets, and in different ICL settings, we show the effectiveness of DNIP. Our key messages are as follows:
-  • We introduce the COBias evaluation metric to assess the pairwise class accuracy bias in language model predictions, and reveal that LLMs of different scales and families consistently exhibit a large COBias score.
-  • We propose Debiasing as Nonlinear Integer Programming (DNIP) to jointly minimize COBias and maximize accuracy.
-An efficient simulated annealing algorithm is adopted to solve the mathematical model.
-  • Experiments show that DNIP achieves the best of both worlds of lower COBias and higher overall accuracy.
+Mitigating class accuracy imbalance is under-explored, but important to facilitate a smoother user experience and reduce safety risks.
+It could be difficult to disentangle all the factors that cause class accuracy imbalance.
+Resolving the issue from the root through pre-training or fine-tuning, leveraging data selection or data augmentation, is neither easy nor cost-effective.
+To overcome the issue, we conceptualize a bias metric, the Contextual Oddity Bias (COBias; Section 2), to quantify class accuracy imbalance from inference-time over-predictions and under-predictions.
+We aim to reduce COBias by directly rectifying the in-context learned LLM output class probabilities.
+We propose DNIP, the Debiasing as Nonlinear Integer Programming model, to search for correction weights leveraging combinatorial optimization, without updating LLM parameters (Section 3).
+These weights adjust class probabilities towards lower COBias and higher overall accuracy.
+On three widely used different-scale LLMs across seven NLP evaluation datasets, and in different ICL settings, we show the effectiveness of DNIP.
+Our key messages are as follows:
+  • We introduce the COBias evaluation metric to quantify the class accuracy imbalance in language model predictions, where LLMs of different scales and families consistently exhibit a large COBias score.
+  • We propose Debiasing as Nonlinear Integer Programming (DNIP) method to find class correction weights to correct ICL output class probabilities, which minimize COBias and maximize accuracy.
+An efficient simulated annealing algorithm is adopted to solve the nonlinear integer programming model.
+  • DNIP achieves the best of both worlds of lower COBias and higher overall accuracy.
 For example, on Llama-2-13B, DNIP significantly reduces COBias (avg.
@@
 2 Over-prediction, Under-prediction, and the Contextual Oddity Bias
-2.1 Over-prediction and Under-prediction in Classification
-[…]
-In the era of Large Language Models (LLMs), over-prediction and under-prediction still persist in LLMs’ predictions, though one may not simply attribute such issues to the black-box training data or the model itself.
-The first obstacle is that the pretraining data’s source distribution for a subgroup of data may not be accessible, making it difficult to directly minimize differences between the source distribution and a downstream target distribution.
-What we normally have is the pretrained model’s output distribution for downstream classes.
-Secondly, the process of prompting an LLM, especially when the prompt is provided with in-context learned demonstrative examples, also brings uncertainty to predictions.
-Ablating the potential biases in prompting is non-trivial.
-More inconveniently, given the relatively large model sizes, it is computationally intensive, expensive, or even infeasible to tweak the LLM training pipeline or conduct LLM fine-tuning to fit debiasing training objectives.
-To see how over-predictions and under-predictions hurt LLM prediction capabilities, we show in Table 2 an example of observed over-predictions and under-predictions in Llama-2-13B.
-We prompt the LLM on the AGNews news classification test set in a one-shot manner, with a single demonstrative example randomly selected from the training set.
-Llama-2-13B greatly over-predicts Business while under-predicting Technology, lowering the LLM’s capability for accurately predicting the Tech class (Tech’s accuracy: 19%).
-In general, there are 2129 predictions in Business but the number of ground-truth instances in Business is 1204; there are only 246 correct Tech predictions but 1240 ground-truth Tech instances.
-Over-/Under-prediction happens in more downstream tasks and larger LLMs (e.g., ChatGPT), as additionally show by examples in Appendix A.
-These surface observations tell us how an LLM could fail in predicting certain classes, suggesting that there is still space for improvements for LLMs to predict every class with equally good accuracy.
-The fact that this problem is difficult to completely eliminate from the source of data, together with the inconveniences in training/fine-tuning debiased LLMs, call for a more actionable inference-time bias metric quantifying the over-/under-predictions and an accompanying debiasing method.
 Pred.
@@
 Prec. 0.91 0.95 0.83 0.84 - -
-Table 2: Significantly reduced over-predictions in class Business and improved test accuracy of the under-predicted class Tech for AGNews after DNIP.
-2.2 Conceptualizing Over-prediction and Under-prediction into the Contextual Oddity Bias
-In Table 2, we observe that Llama-2-13B greatly over-predicts Business while under-predicting Technology (Tech).
-Underneath the surface over-/under-prediction issue, we can see the most significant contributor is the 822 ground-truth instances of Technology, which are wrongly predicted as Business, leading to the model’s low accuracy (19%) in Technology [^1] It simultaneously leads to a fairly low precision score of 55% in Business.
+Table 2: Significantly improved test accuracy of the under-predicted class Tech for AGNews after DNIP correction.
+Over-predictions and under-predictions can hurt LLM prediction capabilities.
+To see how, Table 2 presents an example of observed over-predictions and under-predictions in Llama-2-13B.
+We prompt the LLM on the AGNews news classification [2], and find that Llama-2-13B greatly over-predicts Business while under-predicting Tech, leading to low accuracy in Tech (19%).
+Over-/Under-prediction happens in many downstream tasks and also in larger models such as ChatGPT (Appendix A), suggesting room for improvements so that LLMs can predict all classes with equally good accuracy.
+Conceptualizing Over-prediction and Under-prediction into the Contextual Oddity Bias.
+To introduce the Contextual Oddity Bias, we first elaborate the over-/under-prediction observations.
+In Table 2, the most significant contributor to the under-prediction in Tech (accuracy: 19%) is the 822 ground-truth Tech instances which are wrongly predicted as Business [^1] It simultaneously leads to a fairly low precision score of 55% in Business.
 In fact, the low precision score in Business and the low recall/accuracy score in Technology co-reflect each other.
@@
 Note that by scikit-learn, class recall is the same as class accuracy..
-On that note, we are able to precisely describe the surface over-prediction and under-prediction issue, by decoupling the problem into which output class that a ground-truth class is most wrongly predicted as, suggesting the model’s inference bias towards an output class given a ground-truth class.
-In the AGNews example, Business is what ground-truth Technology is most wrongly predicted as, suggesting the model’s inference bias to predict Technology as Business.
-To quantify this inference bias, we first conceptualize odd classes.
-An odd class is an output class at inference time which holds the most wrong predictions of a ground-truth class $c_{i}$.
-That is, an odd class is always relative to a ground-truth class.
-For $N$ classes and $M$ test instances, let $y$ and $\hat{y}$ denote the ground-truth class and the prediction respectively, and let $u_{c_{i},c_{j}}$ denote the number of ground-truth $c_{i}$ instances that are wrongly predicted as $c_{j}$,
-$u_{c_{i},c_{j}}=\displaystyle\sum\nolimits_{m}\mathbbm{1}\{y_{m}=c_{i},\hat{y}_{m}=c_{j}\},$ (1)
-$odd_class|(c_{i}):=\operatorname*{arg\,max}_{c_{j}}u_{c_{i},c_{j}},$ (2)
-[…]
-For this goal, it could be direct to use the rate of wrong predictions in an odd class as a measurement of its “hurt” to the accuracy of a ground-truth class, but minimizing these rates alone may not adequately mitigating class accuracy imbalances, which is an ultimate goal of reducing over-/under-predictions (More analyses on this measurement are in Appendix B.).
+That is, the output class Business which a ground-truth class Tech is most wrongly predicted as can reflect the model’s inference bias towards an output class for the given ground-truth class.
+To precisely quantify this inference bias, we first conceptualize odd classes.
+An odd class is an output class which holds the most wrong predictions of a ground-truth class.
+That is, there is always an odd class to a ground-truth class.
+For $N$ classes and $M$ instances, let $y$ and $\hat{y}$ denote the ground-truth class and the prediction respectively, let $u_{c_{i},c_{j}}$ denote the number of ground-truth $c_{i}$ instances that are wrongly predicted as $c_{j}$,
+$u_{c_{i},c_{j}}=\displaystyle\sum\nolimits_{m}\mathbbm{1}\{y_{m}=i,\hat{y}_{m}=j\},$ (1)
+$\textrm{odd}_{i}=\textrm{odd\_class}(c_{i}):=\operatorname*{arg\,max}_{c_{j}}u_{c_{i},c_{j}},$ (2)
+where $m=1,...,M,i,j=1,...,N,j\neq i$, $\mathbbm{1}\{\cdot\}=1$ if the condition holds and 0 otherwise.
+For example, odd_class(Tech) = Business.
+Then, we propose to quantify how much an odd class contributes to under-predicting a ground-truth class.
+One may immediately use the ratio of the number of odd class predictions over that of the ground-truth instances, e.g., use 822/1240 to measure the odd Business’s “hurt” to the ground-truth Tech, but minimizing these rates alone may not adequately mitigating class accuracy imbalances (More analyses on this measurement are in Appendix B.).
 A simpler yet better measurement is the absolute accuracy difference between the odd class and the ground-truth class.
-For example, knowing that Business is odd to Tech, the 78% accuracy difference between Business (97%) and Tech (19%) can serve as a measurement of how much Tech is under-predicted.
-The 19% accuracy of Tech also implicitly reflects the many predictions in its odd class Business.
-It should be noted that measuring these accuracy differences sometimes fails to capture the hidden ground-truth-and-odd relationships.
-For instance, World is the odd class to ground-truth Business, but the -12% accuracy difference can not utter this fact.
-Despite this drawback, directly minimizing the accuracy differences will in return significantly reduce predictions in odd classes and improve under-predicted classes’ accuracies.
+For example, the 78% difference between Business (97%) and Tech (19%) quantifies how much Tech is under-predicted.
+The 19% also implicitly reflects the 822 wrong predictions in Business.
+It is noted that these accuracy differences sometimes fail to capture the ground-truth-and-odd relationships.
+For instance, World is the odd class to Business, but the former is 12% (97%-85%) below the latter.
+Despite the drawback, minimizing these differences will in return significantly reduce odd class predictions and improve under-predicted classes’ accuracies.
 The Contextual Oddity Bias.
-To this end, we formulate the definition of Contextual Oddity Bias (COBias).
-First, the absolute value of the accuracy difference between a ground-truth class and its odd class is termed as COBiassingle.
-COBiassingle measures a model’s tendency to over-predict an answer that is caught as odd than the correct answer (or equivalently, a tendency to under-predict the correct answer).
-Notably, for LLMs, an odd answer may be a manifest of surface patterns from prompts or biases in the pretraining data.
-In addition, the odd classes in a same downstream task can vary across different models.
-That’s why we use “contextual” to mean that the oddity of an answer depends on the model or task.
+To this end, we propose Contextual Oddity Bias (COBias).
+First, we propose COBiassingle, the absolute value of the accuracy difference between a ground-truth class and its odd class.
 Formally, given an $N$ classes and $M$ instances, let $x_{m}$ denote an instance whose label is $y_{m}$.
-By few-shot prompting, the model’s output token probability over $N$ classes $c_{1},\dots,c_{N}$ is $\boldsymbol{p}_{m}=(p_{m1},\dots,p_{mN})$.
-$\boldsymbol{p}_{m}$ is normalized over $N$ classes, $\sum_{n=1}^{N}p_{mn}=1$, $m\in\{1,\dots,M\}$.
+After obtaining few-shot prompted softmax probabilities over the vocabulary at the output posistion, we take the probabilities for the class tokens and normalize them to form class probabilities, i.e., $\boldsymbol{p}_{m}=(p_{m1},\dots,p_{mN})$ is the normalized class probabilities for classes $c_{1},\dots,c_{N}$.
 Then the prediction $\hat{y}_{m}$ is $\operatorname*{arg\,max}_{j\in\{1,\dots,N\}}\boldsymbol{p}_{m}$.
 For a given class $c_{i}$, let $A_{\textrm{odd}_{i}}$ and $A_{c_{i}}$ denote the accuracies for the odd class $\textrm{odd}_{i}$ and $c_{i}$ respectively, the COBiassingle metric is computed as:
-$\textrm{odd}_{i}=\verb|odd_class|(c_{i}),$ (3)
 $\textrm{COBias}_{\textrm{single}}=\frac{1}{N}\displaystyle\sum\nolimits_{i}\bigg{|}A_{\textrm{odd}_{i}}-A_{c_{i}}\bigg{|},$ (3)
-where class accuracy $A_{c_{i}}=\frac{1}{|\mathcal{S}_{c_{i}}|}\sum\nolimits_{m\in\mathcal{S}_{c_{i}}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\}$, where $\mathcal{S}_{c_{i}}$ is the set of indices for class $c_{i}$ instances.
+where class accuracy $A_{c_{i}}=\frac{1}{|\mathcal{S}_{c_{i}}|}\sum\nolimits_{m\in\mathcal{S}_{c_{i}}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\}$, $\mathcal{S}_{c_{i}}$ is the set of indices for class $c_{i}$ instances.
 Similarly, $A_{\textrm{odd}_{i}}=\frac{1}{|\mathcal{S}_{\textrm{odd}_{i}}|}\sum\nolimits_{m\in\mathcal{S}_{\textrm{odd}_{i}}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\}$, where $\mathcal{S}_{\textrm{odd}_{i}}$ is the set of indices for class $\textrm{odd}_{i}$ instances.
-Moreover, besides the odd class, which is the most-frequent wrongly predicted class, there could also be a secondly most-, or third most-, to the least-frequent wrongly predicted class.
-For a ground-truth class, there are the odd, the second odd, to the least odd classes.
-For example, for ground-truth Tech (1240 instances), Business is the odd (822), while World is the second odd (156), and Sports is the least odd (27).
-Then, we extend the accuracy difference between the single odd class and a ground-truth class to every pair of classes, and absolute value can be taken.
+COBiassingle measures a model’s tendency to under-predict the true answer and mostly mistake it as an odd answer.
+The odd classes in a same downstream task can vary across different models.
+That’s why we use “contextual” to mean that the oddity of an answer depends on the model or task.
+Beyond the single odd classes, there could also be a secondly most-, or third most-, to the least-frequent wrongly predicted class, i.e., the odd, the second odd, to the least odd classes.
+For example, for ground-truth Tech, Business is the odd, while World is the second odd, and Sports is the least odd.
+We extend $\textrm{COBias}_{\textrm{single}}$ to every pair of classes.
 This arrives at the ultimate the COBias metric as follows.
 $\textrm{COBias}={N\choose 2}^{-1}\displaystyle\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\bigg{|}A_{c_{i}}-A_{c_{j}}\bigg{|}.$ (4)
-This COBias metric directly reflects the inter-class imbalance in accuracy.
-Indeed, the absolute value of accuracy differences of some pairs in Equation 5 may mask the most frequent wrongly predicted class (the odd class).
-Nonetheless, it reflects a well-rounded measurement of all pairs of classes, whether involving odd classes or not, so it is a powerful metric when we strongly demand minimal class accuracy imbalance.
-In a nutshell, the debiasing problem is minimizing COBias (Eq.
-5) while maintaining or increasing the overall inference accuracy.
-This post-hoc bias evaluation metric will enable easy-to-operate inference-time/post-hoc output likelihood optimization, lifting the heavy computations required for training-wise data selection or model parameter updates.
-3 Improving LLM Outputs by Debiasing as Nonlinear Integer Programming
+COBias directly reflects the inter-class imbalance in accuracy.
+Although the absolute value of accuracy differences of some pairs in Equation 4 may mask the odd class, it reflects a well-rounded measurement and a powerful metric when minimal class accuracy imbalance is demanded.
 Figure 1: An overview of the DNIP method.
@@
 ($0\rightarrow 66\%$).
-In this work, we approach the debiasing problem as a nonlinear integer programming (NIP) problem.
-The objective is to optimize correction weights to rectify the biased, class probabilities, by directly minimizing the COBias metric while maximizing two accuracy-related metrics.
-The weights are optimized on a set of labeled examples, and are simply plugged in during test time.
-An overview of the proposed DNIP method is shown in Figure 1.
+3 Improving LLM Outputs by Debiasing as Nonlinear Integer Programming
+Essentially, we aim to rectify biased few-shot prompted LLM class probabilities with correction weights, so that corrected class probabilities should lead to less class accuracy imbalance.
+With this goal, we propose an inference-time nonlinear integer programming (NIP) selection model to search for optimal correction weights from a weight scale that directly minimizes COBias while maximizing overall accuracy.
+Therefore, the correction weights are optimized by the NIP model on a set of given LLM output class probabilities with ground-truth labels.
+For a test instance, after obtaining its class probabilities, the learned correction weights are applied to correct them.
+An overview of the proposed Debiasing as Nonlinear Integer Programming (DNIP) method is shown in Figure 1.
 3.1 The DNIP Objective
-To debias $\boldsymbol{p}_{m}$, we introduce a discrete $K$-point correction scale $\boldsymbol{\omega}$ ranging from 0 to 1, to re-weight class probability.
-Let $\boldsymbol{\omega}=(\omega_{1},\dots,\omega_{K})$ denote the parameter for the discrete weights in the correction scale, where $\omega_{k}=k/K$.
-An example $10$-point scale is given by $\boldsymbol{\omega}=(0.1,\dots,1.0)$.
-For each class probability in $\boldsymbol{p}_{m}$, it multiplies by a correction weight chosen from $\boldsymbol{\omega}$ to obtain a debiased class probability.
-Finding optimal correction weights is a combinatorial optimization problem: for $N$-way class probability $\boldsymbol{p}_{m}$, each class selects a weight from the $K$-dimensional correction scale.
-To find the optimal combination, the brute force strategy is an enumerated search, where the number of weight combinations to explore is $K^{N}$ for an example; for $M$ examples, it results in a computational complexity of $\mathcal{O}(K^{N}\cdot M)$, which can not run in polynomial time, calling for smarter methods.
-Instead, we model the debiasing problem as a nonlinear integer program, and search for an optimal set of class correction weights with simulated annealing.
-Without loss of generality, we boil down the problem to selecting optimal weight indices towards minimizing the pairwise class accuracy imbalance, while maximizing the overall accuracy and the pointwise mutual information between class labels and predictions.
-Formally, we define an integer selection variable $\xi=(\xi_{1},\dots,\xi_{N})$ that selects a weight index $k$ for the $n$-th class probability in $\boldsymbol{p}_{m}$, namely:
-$\xi_{n}=k\quad\text{if the $n$-th class selects }\omega_{k}.\$ (6)
-The DNIP objective function and constraints are formally given by:
-$\min\quad$ $Z(\xi)=\frac{1}{M}\displaystyle\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}+\beta{N\choose 2}^{-1}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\bigg{|}A_{i}-A_{j}\bigg{|}-\tau\displaystyle\sum_{j=1}^{N}\textrm{PMI}_{j}$ (7)
+We first introduce a discrete $K$-point correction weight scale $\boldsymbol{\omega}$ to select correction weights from.
+We denote $\boldsymbol{\omega}=(\omega_{1},\dots,\omega_{K})$, and $\omega_{k}=k/K$ is a value from the scale.
+For example, a $10$-point scale renders $\boldsymbol{\omega}=(0.1,\dots,1.0)$.
+To debias the $N$-dimensional $\boldsymbol{p}_{m}$, for each class $j$, $\boldsymbol{p}_{mj}$ multiplies a correction weight selected from $\boldsymbol{\omega}$ to obtain a corrected class probability.
+Therefore, finding optimal correction weights for all $N$ classes becomes a combinatorial optimization problem: each class selects a weight from the $K$-point scale, so we need to find the optimal combination of the selections.
+To this end, we formulate the following correction weight selection model towards minimizing pairwise class accuracy differences and maximizing overall accuracy.
+Formally, we define integer selection variables $\xi=(\xi_{1},\dots,\xi_{N})$ that selects a weight index $k$ for the $n$-th class probability in $\boldsymbol{p}_{m}$, namely, $\xi_{n}=k$ if the $n$-th class selects $\omega_{k}$.
+As a result, the DNIP model can be given by:
+$\min\quad$ $Z(\xi)=\frac{1}{M}\displaystyle\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}+$
+$\beta{N\choose 2}^{-1}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}\bigg{|}A_{c_{i}}-A_{c_{j}}\bigg{|}-\tau\displaystyle\sum_{j=1}^{N}\textrm{PMI}_{j}$ (5)
 s.t. $\hat{y}_{m}=\operatorname*{arg\,max}_{n\in\{1,\dots,N\}}\omega_{\xi_{n}}p_{mn},\qquad m=1,\dots,M$ (6)
-$A_{i}=\frac{1}{|\mathcal{C}_{i}|}\sum_{m\in\mathcal{C}_{i}}\mathbbm{1}\{\hat{y}_{m}=y_{m}\},\qquad i=1,\dots,N$ (9)
-$\textrm{PMI}_{j}=\textrm{PMI}(\hat{S}^{j},S^{j})=\log\frac{f(\hat{S}^{j},S^{j})}{f(\hat{S}^{j})f(S^{j})},\qquad j=1,\dots,N$ (10)
+$\textrm{PMI}_{j}=\textrm{PMI}(\hat{S}^{j},S^{j})=\log\frac{f(\hat{S}^{j},S^{j})}{f(\hat{S}^{j})f(S^{j})},$
+$j=1,\dots,N$ (7)
 $\xi_{n}\in\{1,\dots,K\},\qquad n=1,\dots,N$ (8)
-where $\beta$ and $\tau$ are parameters that control the bias and the negative PMI terms respectively, $\hat{S}^{j}$ and $S^{j}$ denote the examples of prediction $j$ and true label $j$ respectively.
-In details, $f(\hat{S}^{j})$ is the ratio between the number of examples with prediction $j$ and the total number of examples, similarly, $f({S}^{j})$ is the ratio between the number of examples with class label $j$ and the total number of examples, $f(\hat{S}^{j},S^{j})$ is the ratio between the number of correct predictions of class $j$ and the total number of examples.
-Therefore, $\textrm{PMI}_{j}=\log\big{(}\frac{\mathrm{n}(\hat{S}^{j},S^{j})/N}{\mathrm{n}(\hat{S}^{j})/N\cdot\mathrm{n}(S^{j})/N}\big{)}$, where $\mathrm{n}(\cdot)$ is the count.
-In actual computations, we apply add-$u$ smoothing to smooth out zeros in the numerator and denominator.
-The value of the smoothing coefficient $u$ is selected along other parameters on a development set.
-The debiased class probabilities and prediction are:
-$\boldsymbol{p}^{\ast}_{m}$ $=(\omega_{\xi_{1}}p_{m1},\dots,\omega_{\xi_{N}}p_{mN})$ (12)
-$y^{\ast}_{m}$ $=\operatorname*{arg\,max}_{n\in\{1,\dots,N\}}\boldsymbol{p}^{\ast}_{m}$ (13)
-The intuition is to encourage the adjusted predictions to have lower inter-class bias while improving overall accuracy.
-As the second term COBias has minimized the inter-class accuracy bias, the first term error rate brings the class prediction closer to the actual class, and the third term negative PMI further enforces the class prediction to be more confident when it is close to the actual class.
-The PMI term can also be seen as a constraint to penalize each individual class.
+where $\beta$ and $\tau$ are parameters that control the bias and the PMI terms respectively.
+The PMI term is introduced to maximize the pointwise mutual information between ground-truth class $j$ instances and predicted class $j$, which can be seen as a confidence level, or a constraint to penalize each individual class.
+$\hat{S}^{j}$ and $S^{j}$ denote the instances of prediction $j$ and true label $j$ respectively, $f(\hat{S}^{j})$ is the ratio between the number of instances with prediction $j$ and the total number of instances, similarly, $f({S}^{j})$ is the ratio between the number of instances labeled $j$ and the total number of instances, $f(\hat{S}^{j},S^{j})$ is the ratio between the number of correct predictions of class $j$ and the total number of instances.
+In actual computations, we apply add-$\mu$ smoothing to the respective ratios in the numerator and denominator to smooth out zeros.
+The smoothing coefficient $\mu$ is selected along other parameters on a development set.
+Intuitively, the first term directly minimizes error rates, the second term COBias directly minimize the inter-class accuracy differences, and the PMI term, is employed to further encourage predictions within a class to be close to the actual class.
+The corrected class probabilities and prediction are:
+$\boldsymbol{p}^{\ast}_{m}$ $=(\omega_{\xi^{\ast}_{1}}p_{m1},\dots,\omega_{\xi^{\ast}_{N}}p_{mN})$ (9)
+$\hat{y}^{\ast}_{m}$ $=\operatorname*{arg\,max}_{n\in\{1,\dots,N\}}\boldsymbol{p}^{\ast}_{m}$ (10)
 3.2 Solving DNIP with Simulated Annealing
@@
 16:return $\xi^{\ast},y^{\ast}$
-We solve DNIP by simulated annealing [6, 7].
-Simulated annealing (SA) is a Metropolis-Hastings sampling algorithm, which is widely adopted to solve optimization problems with its versatility in dealing with highly nonlinear models, noisy data, and many constraints [8].
-The DNIP mathematical model is such a case, pointing to the need of SA.
-We especially take advantages of SA algorithms’ searching ability - escaping from local optima and finding global optima after reaching local optima - to find optimal correction weights.
-In details, the SA strategy consists of searching the solution space starting from a randomly initialized $\xi$ and then generating a new one by perturbing it.
-The new solution is either accepted or rejected by the Metropolis criterion after evaluating the objectives.
+Because the NIP objective contains discontinuous functions and is non-differentiable, we solve DNIP by simulated annealing [4, 5].
+Simulated annealing (SA) is a Metropolis-Hastings sampling algorithm, which is widely adopted to solve optimization problems with its versatility in dealing with nonlinear models and many constraints [6].
+The DNIP mathematical model is nonlinear, so SA is suitable.
+The main advantage of SA is escaping from local optima and finding global optima after reaching local optima, which is helpful in finding optimal correction weights.
+In details, the SA algorithm consists of searching the solution space starting from a randomly initialized $\xi$ and generating a new one by perturbing it.
+The new solution is either accepted or rejected by the Metropolis criterion after evaluating the DNIP objective.
 Algorithm 1 describes the searching for optimal $\xi$ using SA.
 The outer loop performs the cooling/annealing process.
-In this work, a geometric decay with $\alpha=0.95$ and initial temperature of 200,000 is set as cooling schedule.
-The inner loop simulates the thermal equilibrium reaching process at a temperature, and its criterion ensures that either the number of generated or accepted solutions at a temperature is above a threshold [9].
+In this work, the cooling schedule is set with a geometric decay with $\alpha=0.95$ and initial temperature of 200,000.
+The inner loop simulates the thermal equilibrium reaching process at a temperature, and the criterion ensures that either the number of generated or accepted solutions at a temperature is above a threshold [7].
 We sample a new $\xi$ from the neighborhood by randomly substitute a selection in $\xi$.
-The computational complexity for the SA algorithm is $\mathcal{O}(NK)$.
+The computational complexity for SA is $\mathcal{O}(NK)$.
 Please refer to Appendix C for derivations.
@@
 Evaluation Tasks.
-The proposed method is evaluated on a diverse set of binary and multi-class NLP classification tasks, across general domains and biomedical domains.
-The five general-domain evaluation datasets are 4-way news topic classification AGNews [10], 14-way ontology classification DBpedia [11], 5-way sentiment classification SST-5 [12], 6-way retrieval question classification TREC [13, 5], binary entailment recognition RTE [14]; the two domain-specific datasets from BLURB [15] are 5-way biomedical relation extraction DDI [16], and 3-way biomedical question answering PubMedQA [17].
+We evaluate on a diverse set of binary and multi-class NLP classification tasks, across general domains and biomedical domains.
+The five general-domain evaluation datasets are 4-way news topic classification AGNews, 14-way ontology classification DBpedia [8], 5-way sentiment classification SST-5 [9], 6-way retrieval question classification TREC, binary entailment recognition RTE [10]; the two domain-specific datasets from BLURB [11] are 5-way biomedical relation extraction DDI [12], and 3-way biomedical question answering PubMedQA [13].
 The selected biomedical tasks are of practical uses.
@@
 Each dataset comes with a training and a test set.
-We make use of the default training set for DNIP optimization, except for AGNews, DBpedia, and DDI, for which we randomly select 10,000 training examples since the datasets are quite large.
-We then split the training examples into 95%, 5% optimization and development sets.
-We evaluate on the default test set except for AGNews, for which we evaluate on 5,000 randomly selected test examples.
-Evaluation metrics are COBias and accuracy.
+We take the full or subset of the training set and obtain ICL class probabilities with ground-truth labels as the optimization set for DNIP optimization.
+Evaluations are performed on each test set against COBias and overall accuracy.
+Optimization and dev splits can be found in Appendix D.
 Models and Experimental Configuration.
-We evaluate three open-source LLMs across varied scales and model families, including GPT-2-XL (1.5B parameters), Llama-2-7B (7B parameters), and Llama-2-13B (13B parameters).
-For main results, we use 1-shot prompting and obtain the ICL output class softmax probabilities.
-To standardize calculations, we normalize the softmax probabilities over all classes.
-For DNIP, we tune parameters of $\beta,\tau,K,u$ on the development set.
-To account for variance caused by different demonstrations, for each model and dataset, we perform three runs of prompting with different 1-shot demonstrations and obtain three sets of initial probabilities.
-For both ICL and DNIP, we report the average test accuracy and COBias over the three runs.
-The demonstration is randomly selected from training examples.
-All prompting is done on an NVIDIA A100 GPU.
-Simulated annealing executes on CPU (execution time depends on the optimization set size, 3 to 30 minutes on AMD EPYC 7742 CPU or slightly longer on local laptop CPU).
+We evaluate three open-source LLMs across varied scales and families, including GPT-2-XL (1.5B parameters), Llama-2-7B, and Llama-2-13B.
+We do few-shot prompting on an NVIDIA A100 GPU, main results use 1-shot prompting, and we also perform more few-shot analyses.
+For DNIP, $\beta,\tau,K,\mu$ are tuned on the development set.
+To account for variance caused by demonstrations in prompts, we perform three runs for each dataset with different 1-shot demonstrations (randomly selected from the optimization set).
+Average scores over three runs are reported.
+SA executes on CPU.
 Model Eval. Metric AGNews DBpedia SST-5 TREC RTE DDI PubMedaQA Avg.
@@
 4.2 Main Results
-The average test accuracy (in black) and COBias (in blue) with standard deviation over three runs are shown in Table 3.
-In all three LLMs, test COBias by ICL is large, with GPT-2-XL exhibiting the largest COBias (50.3%), showing that class accuracy imbalance exists in LLMs of different sizes and families.
-With DNIP, test COBias is significantly reduced, by an absolute 29%, 23%, and 26% for the three models respectively; meanwhile, there is an 16%, 12%, and 10% absolute improvement in test accuracy respectively.
-It is worth noting that for tasks whose ICL test accuracy is already in higher 80s, e.g., DBpedia, DNIP can reduce COBias while further improving the accuracy to 90s.
-In addition, we observe a very large ICL COBias at 82.4% in GPT-2-XL for RTE, reflecting severe over-predictions in class True.
-After applying DNIP, the COBias greatly reduces to 7.1%.
-That is, DNIP obtains over 75% absolute reduction in COBias for RTE, demonstrating DNIP’s capability for balancing very imbalanced class accuracies without hurting the overall accuracy much.
-Moreover, on the two biomedical tasks, DNIP brings an average absolute reduction of 31% and 13% in COBias, and an average absolute increase of 34% and 11% in accuracy, showing its effectiveness for domain-specific tasks.
+Averaged test accuracy (in black) and COBias (in blue) with standard deviation over three runs are shown in Table 3.
+Seen from the ICL COBias scores, class accuracy imbalance exists in LLMs of different sizes and families.
+With DNIP, test COBias is significantly reduced by an absolute 29%, 23%, and 26% for the three models respectively; test accuracy is improved by 16%, 12%, and 10% respectively.
+For tasks whose ICL test accuracy is already in higher 80s, e.g., DBpedia, DNIP can reduce COBias while further improving the accuracy to 90s.
+In addition, ICL COBias is very large (82.4%) on GPT-2-XL for RTE, reflecting severe under-predictions in class False.
+After applying corrections, its COBias greatly reduces to 7.1%, demonstrating DNIP’s capability for balancing very imbalanced class accuracies without hurting the overall accuracy much.
+DNIP is also effective for biomedical tasks, bringing an average absolute reduction of 31% and 13% in COBias, and an average absolute increase of 34% and 11% in accuracy.
+Dataset Time/s Variable Size
+AGNews 2278 120
+DBpedia 2584 420
+SST-5 1499 150
+TREC 1274 180
+RTE 518 60
+DDI 2092 150
+PubMedQA 203 90
+Table 4: Optimization time (seconds).
 Moreover, the 13B Llama model can suffer even higher COBias than its 7B counterpart on biomedical tasks, further showing that class accuracy imbalance does not go away as model sizes increase, and DNIP is effective to balance class accuracies for LLMs of all sizes.
-Back to the AGNews example in Table 2, we present the DNIP debiased results in Table 2.
-Over-predictions in Business are greatly reduced, and the accuracy of Tech is significantly improved from 19% to 84%.
-These results show that DNIP strongly mitigates COBias and greatly improves the accuracy of lower-performing classes, while boosting overall accuracy.
+Furthermore, we present the confusion matrix of DNIP debiased AGNews results in Table 2.
+Under-predictions in Tech are greatly reduced, with its accuracy significantly improved from 19% to 84%.
+Additionally, DNIP requires only a small optimization set of 10 examples for a fair improvement in both COBias and accuracy, and it exhibits a further, emergent COBias reduction at several thousands optimization examples.
+Figure 2: Objective function and temperature curves.
+We discuss performance improvements vs. the number of optimization examples needed in Appendix E.
 Computational Time.
-Regarding computational time needed to perform DNIP optimization, on all datasets, the optimization is finished in the scale of minutes.
-Though computational time increases with more classes and a more fine-grained weight scale, it should be noted that the time elapsed in optimization is dozens of minutes even for tasks of more than 10 classes, and a weight scale of 30 is good enough for most of the datasets experimented in this paper.
-This demonstrates the practicality of DNIP in the computational time needed for performing optimization.
+For all datasets, optimization finishes in minutes.
+Though the optimization time varies across variable size ($N\cdot K$) and optimization set size, it finishes in dozens of minutes even for DBpedia (14 classes; Figure 2).
+A weight scale of 30 is good enough for most datasets evaluated.
+Optimization time using weight scale 30, $\beta=2.7,\tau=0.2$ is shown in Table 4, demonstrating the practicality of DNIP as an inference-time debiasing method without incurring much costs.
+4.3 More ICL Settings
 Prompting Method Metric AGNews DBpedia SST-5 TREC RTE DDI PubMedQA Avg.
@@
 Acc $88.5_{0.6}$ $95.8_{0.7}$ $52.9_{2.5}$ $76.6_{8.7}$ $75.8_{4.5}$ $54.1_{5.1}$ $59.9_{1.4}$ 71.9
-5-shot DNIP COBias $7.0_{0.9}$ $5.7_{1.1}$ $15.8_{12.6}$ $14.4_{7.5}$ $3.1_{1.9}$ $16.5_{7.6}$ $9.6_{4.6}$ 10.3
+DNIP COBias $7.0_{0.9}$ $5.7_{1.1}$ $15.8_{12.6}$ $14.4_{7.5}$ $3.1_{1.9}$ $16.5_{7.6}$ $9.6_{4.6}$ 10.3
 Acc $83.5_{1.5}$ $95.2_{1.2}$ $50.3_{2.3}$ $67.0_{12.7}$ $75.0_{0.8}$ $9.7_{1.0}$ $52.3_{5.3}$ 61.9
@@
 Acc $88.7_{0.5}$ $96.6_{0.5}$ $51.3_{1.0}$ $82.7_{1.4}$ $76.7_{3.7}$ $43.6_{6.1}$ $55.3_{2.6}$ 70.7
-k-shot (1 demon. per class) DNIP COBias $7.3_{0.5}$ $4.3_{0.7}$ $2.8_{1.5}$ $12.1_{5.5}$ $5.0_{3.3}$ $12.5_{4.3}$ $8.7_{1.2}$ 7.5
+DNIP COBias $7.3_{0.5}$ $4.3_{0.7}$ $2.8_{1.5}$ $12.1_{5.5}$ $5.0_{3.3}$ $12.5_{4.3}$ $8.7_{1.2}$ 7.5
 Table 5: DNIP results in more ICL settings. Average score over three runs are reported.
-4.3 More ICL Settings
-We prompted Llama-2-13B in two additional settings: 5-shot, where each demonstration is randomly selected as in the 1-shot case; k-shot, where k is the number of classes and each class is represented by a demonstration example in the prompt.
-Table 4 presents results on seven benchmark datasets in both settings.
-For example, the relative average COBias reduction (comparing DNIP to ICL) is 65% and 74% for 1-shot and 5-shot cases respectively, and the relative average accuracy increase is 16% and 11% for 1-shot and 5-shot cases respectively, demonstrating DNIP’s capabilities with more shots.
-In addition, prompt design does contribute to different output probabilities, which could be helpful to provide different starting points for DNIP, so DNIP could optimize for better solutions.
-However, prompt engineering alone may not be most effective to solve the COBias issue.
-For example, compared to 1-shot prompting (Table 3), increasing the number of shots does not significantly further reduce COBias or boost accuracy; using more diverse demonstrations (k-shot) may help with COBias, but does not gain much higher accuracy than 1-shot.
-Therefore, using more shots does not help reduce COBias much, and DNIP significantly reduces COBias and improves accuracy in both 5-shot and k-shot settings, further showcasing its effectiveness.
-4.4 DNIP Significantly Reduces Over-predictions, and Balances Class Accuracies across Varied Classification Tasks, as many as 14 Classes
-Figure 2: Pairwise class accuracy differences on test sets before and after applying DNIP; the pinker the higher class accuracy difference, the greener the lower; average accuracy over three runs are used.
-Heatmaps after DNIP show nearly 0 accuracy differences for many class pairs.
+We prompt Llama-2-13B in two additional settings: 5-shot, where each demonstration is randomly selected; k-shot, where k is the number of classes and each class is represented by a demonstration, randomly selected.
+Table 5 shows that, for example, the relative average COBias reduction by DNIP is 65% and 74% for 1-shot (Table 3) and 5-shot respectively, with relative average accuracy increase of 16% and 11%.
+The rationale is, demonstrations contribute to variance in outputs, but prompt engineering alone does not mitigate much the imbalance.
+DNIP can correct the outputs no matter what different starting points are given.
 Figure 3: Test confusion matrix for DBpedia before and after applying DNIP. DNIP greatly improves ICL predictions for Nat., which is the most under-predicted class by ICL.
-We visualize pairwise class accuracy differences before and after applying DNIP to 1-shot, Llama-2-13B ICL results in Figure 2.
-In each heatmap, row $i$, column $j$ cell represents the absolute difference between the test accuracies of two classes in row $i,j$ , i.e., $|A_{i}-A_{j}|$.
+4.4 DNIP Significantly Balances Class Accuracies for Classification Tasks with as Many as 14 Classes
+We further investigate DBpedia.
+By ICL, class Nat.
+exhibits the largest accuracy differences to others, and it is the most under-predicted class, whose odd class is Trans.
+and second odd class is Bldg., shown in Figure 3 (using seed 0 run).
+DNIP greatly reduces Trans.
+and Bldg.
+predictions for Nat..
+When averaged over all three runs, test accuracy for Nat.
+improves from 44% to 89%, and that for Trans.
+increases from 78% to 87%, demonstrating DNIP’s competence in reducing class accuracy differences.
+Figure 4: Pairwise class accuracy differences on test sets before and after DNIP; the pinker the higher difference, the greener the lower; averaged over three runs.
+DNIP achieves nearly 0 accuracy differences for many class pairs.
+We also visualize pairwise class accuracy differences before and after applying DNIP to 1-shot, Llama-2-13B ICL results in Figure 4.
+In each heatmap, row $i$, column $j$ cell represents the absolute difference between the test accuracies of two classes in row $i,j$ , i.e., $|A_{c_{i}}-A_{c_{j}}|$.
 DNIP brings the accuracy differences down to close to zero for AGNews, DDI, DBpedia, and RTE.
 The lighter green heatmaps in the bottom row show that DNIP consistently reduces COBias across different NLP classification tasks.
-Relieving Odd Classes for 14-Way Ontology Classification.
-We further investigate the 14-way classification task DBpedia.
-Before debiasing, the Nat.
-class exhibits the largest class accuracy difference to other classes.
-In fact, Nat.
-is the most under-predicted class, whose odd class is Trans.
-and second odd class is Bldg., as shown by the confusion matrix in Figure 3 (based on seed 0, 1-shot ICL run).
-DNIP greatly reduces false negatives for Nat., i.e., reduces wrong Trans.
-and Bldg.
-predictions for Nat..
-Moreover, on an average over all three runs, the test accuracy for Nat.
-improves from 44% to 89%, and that for Trans.
-also increases from 78% to 87%, demonstrating that DNIP is effective in reducing class accuracy differences for tasks of as many as 14 classes.
-Furthermore, we present more heatmaps for smaller LLMs including Llama-2-7B and GPT-2-XL in Appendix D.
-4.5 Optimizing COBiassingle
-Using COBiassingle (Eq. 4) for DNIP optimization will focus on odd classes only. Figure 4 shows ablation results when replacing the COBias term in the DNIP objective with COBiassingle.
-$\min\quad Z(\xi)=\frac{1}{M}\displaystyle\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}+\beta\frac{1}{N}\displaystyle\sum\nolimits_{i}\bigg{|}A_{\textrm{odd}_{i}}-A_{c_{i}}\bigg{|}-\tau\displaystyle\sum_{j=1}^{N}\textrm{PMI}_{j}$ (14)
-Figure 4: Test COBiassingle, COBias, and overall accuracy scores (from left to right).
-Blue dots represent the respective evaluation metric score of the original 1-shot ICL Llama-2-13B outputs of the seven benchmark datasets; green dots are debiased results optimized by the COBiassingle DNIP objective (Eq.
-14); red dots are debiased results optimized by the COBias DNIP objective.
-Comparing DNIP with COBiassingle and DNIP with COBias, their performances are close, but DNIP with COBias achieves higher overall test accuracy and slightly better bias mitigation performances.
-In other words, DNIP with COBiassingle can significantly minimize accuracy differences between odd classes and ground-truth classes, but DNIP with COBias is even more effective when it comes to balancing all pairs of classes, not limited to the odd classes.
-In summary, optimizing either COBiassingle or COBias is an effective inference-time measure to reduce over-predictions in raw ICL outputs and boost accuracy for under-predicted classes.
-4.6 Accuracy-related Objectives vs. the COBias Objective
+Furthermore, we present more heatmaps for smaller LLMs including Llama-2-7B and GPT-2-XL in Appendix F.
+Figure 5: Ablation analysis of the objective function, exemplified with Llama-2-13B; test scores averaged over 7 datasets are used for the rightmost plot.
+4.5 Accuracy-related Objectives vs. COBias Objective
 We show that the integration of accuracy-related objectives and the COBias objective are indispensable in achieving the best of both worlds in Figure 5.
-DNIP objective function $z$ consists of three terms: the error rate term $z_{1}=\frac{1}{M}\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}$; the COBias term $z_{2}={N\choose 2}^{-1}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}|A_{i}-A_{j}|$; and the PMI term $z_{3}=\sum_{j=1}^{N}\textrm{PMI}_{j}$; the accuracy-related terms are $z_{1}$ and $z_{3}$.
+DNIP objective function $z$ consists of three terms: the error rate term $z_{1}=\frac{1}{M}\sum_{m=1}^{M}\mathbbm{1}\{\hat{y}_{m}\neq y_{m}\}$; the COBias term $z_{2}={N\choose 2}^{-1}\sum_{i=1}^{N-1}\sum_{j=i+1}^{N}|A_{c_{i}}-A_{c_{j}}|$; and the PMI term $z_{3}=\sum_{j=1}^{N}\textrm{PMI}_{j}$; the accuracy-related terms are $z_{1}$ and $z_{3}$.
 We ablate all ablation objective functions to perform DNIP: $z_{1},z_{2},z_{3},z_{1}+\beta z_{2},z_{1}-\tau z_{3},\beta z_{2}-\tau z_{3},z_{1}+\beta z_{2}-\tau z_{3}$.
-Ablations on GPT-2-XL and Llama-2-7B show similar findings (Appendix E).
+Ablations on GPT-2-XL and Llama-2-7B show similar findings (Appendix G).
 To make fair comparisons, we fix $\beta=2.7,\tau=0.2$ and weight scale $K=30$, which may not result in the same scores reported in Table 3.
@@
 (2) Using $z_{1}-\tau z_{3}$ also obtains a higher/similar accuracy but a higher COBias.
-(3) Using $z_{2}$ solely focuses on mitigating COBias and is straightforwardly effective in reducing COBias.
-However, the shortcoming is obvious: a lower accuracy.
-(4) Using $z_{1}+\beta z_{2}$ or $\beta z_{2}-\tau z_{3}$ is promising with a similar/lower accuracy and a slightly lower COBias.
-We note that accuracy-related objectives alone increase accuracy by over 16% ($75\%-59\%$), which is greater than using the objective combining COBias, suggesting DNIP is effective for those who emphasize accuracy improvement.
-Figure 5: Ablation analysis of the objective function, exemplified with Llama-2-13B; test scores averaged over 7 datasets are used for the rightmost plot.
-4.7 Performance Improvements vs. Number of Optimization Examples
-We evaluate DNIP with varying optimization examples on Llama-2-13B.
-We find that DNIP requires only a small optimization set of 10 examples for a fair improvement in both COBias and accuracy, and it exhibits a further, emergent COBias reduction at several thousands optimization examples.
-Figure 6: Relative test COBias reduction; higher is better.
-Test COBias reduction reaches 9% to 72% at 10 optimization examples, and continues to grow with more optimization examples.
-Figure 6 shows the relative test COBias reduction over ICL results.
-At 10 examples, DNIP can achieve a relative COBias reduction ranging from 9% to 72%.
-We observe an overall trend of COBias reduction except for DDI and PubMedQA, who have a spike at 100 examples, but DDI rises again after 500 examples while PubMedQA levels off (PubMedQA full optimization size: 950).
-An explanation is that DNIP prioritizes COBias reduction more than accuracy increase at fewer examples; at more examples, the accuracy-related objectives show more influence.
-It also indicates that domain-specific datasets are more sensitive to the trade-off.
-We further notice a surge in reductions from 1,000 to full for DDI, SST-5, and DBpedia (full optimization sizes: 9,500, 8,116, 9,500), suggesting that around 8,000 examples are needed for a further emergent reduction - PubMedQA may see further COBias reduction if more optimization examples are available.
-Test accuracy is improved with only 10 optimization examples, and plateaus at around 500 examples.
-Figure 7 shows that, at 10 examples, the relative increase in accuracy ranges from 0.2% to 11% except DDI, which is 160%.
-At 100 examples, the increase ranges from 3% to 208% (DDI, $7.2\%\rightarrow 18.7\%$).
-The increase continues to grow and stabilizes at around 500 examples.
-At 1,000 examples, test accuracy for DDI rises to 52%, which is more than 6 times of the ICL accuracy, showing that DNIP is effective for this challenging biomedical NLP task.
-Figure 7: Relative test accuracy increase; higher is better.
-Moreover, we observe for RTE a declining trend of accuracy increase, which matches its growing trend of COBias reduction, reflecting the trade-off between the two objectives.
-We also note a slight decrease in accuracy improvement for SST-5 and DDI at full size, which matches their sudden increase in COBias reduction, verifying that DNIP has emergent COBias reduction abilities with thousands of optimization examples.
-Overall, DNIP improves both accuracy and COBias at 10 examples, and achieves acceptable COBias reduction and accuracy improvement at 100 examples, with test COBias reduced by 57% and test accuracy improved.
-For detailed scores, please refer to Appendix F.
+(3) Using $z_{2}$ solely focuses on mitigating COBias and is straightforwardly effective in reducing COBias, but the shortcoming is a lower accuracy.
+(4) Using $z_{1}+\beta z_{2}$ or $\beta z_{2}-\tau z_{3}$ is promising with a similar/lower accuracy and a lower COBias.
+Accuracy-related objectives alone increase accuracy by over 16% ($75\%-59\%$), suggesting DNIP is effective for those who emphasize accuracy improvement.
+Figure 6: Comparisons to more methods.
+4.6 Optimizing COBiassingle
+Using COBiassingle (Eq.
+3) for DNIP optimization will focus on odd classes only.
+Figure 7 shows ablation results when replacing the COBias term in the DNIP objective with COBiassingle.
+Blue dots represent the respective evaluation metric score of the original 1-shot ICL Llama-2-13B outputs of the seven benchmark datasets; green dots are debiased results optimized by the COBiassingle DNIP objective; red dots are results optimized by the COBias DNIP objective.
+Figure 7: Test COBiassingle, COBias, and overall accuracy scores (from left to right).
+Comparing DNIP with COBiassingle and DNIP with COBias, their performances are close, but DNIP with COBias achieves higher overall test accuracy and slightly better bias mitigation performances.
+In other words, DNIP with COBiassingle can significantly minimize accuracy differences between odd classes and ground-truth classes, but DNIP with COBias is even more effective when it comes to balancing all pairs of classes, not limited to the odd classes.
 5 Related Work
-LLMs may fail in various ways.
-Analysis towards understanding LLM failure modes has gained remarkable insights on how they fail and how to identify/avoid failures.
-For example, understanding hallucinations in LLMs facilitates fact-checking evaluation metrics for LLMs.[18, 19, 20].
-Jones and Steinhardt [21] reveals failure modes of LLMs which resemble cognitive biases and measures the failures on GPT-3.5 and Codex.
+Class accuracy imbalance is one of the potential failures of LLMs.
+Analysis towards understanding LLM failure modes has gained remarkable insights.
+For example, understanding hallucinations in LLMs facilitates fact-checking metrics for LLMs.[14, 15, 16].
+Jones and Steinhardt [17] reveals failure modes of LLMs which resemble cognitive biases.
 Lu et al.
-[22] identify ICL order sensitivity that result in LLM prompting failures.
-Turpin et al.
-[23] shows that LLM reasoning abilities are affected by adding biasing features to the input.
-Along this line, we tackle a common LLM failure of the class accuracy difference, via nonlinear integer programming optimization on the evaluation metric scores.
-The inference-time class accuracy bias has multi-faceted causes.
-It could also be attributed to pre-training data, such as predilections for common tokens in the pretraining corpora [24].
-It could be a direct outcome of LLMs’ liking for specific patterns or answers, leading to highly imbalanced predicted class distributions [25].
-Other causes include choices of prompt templates [26, 27, 28], demonstrations [29], and prompt orders [22, 23].
-Unintended LLM behaviors may also exacerbate the bias, such as user-following sycophancy [30, 31, 32] and prompt toning effects [21, 33, 34, 35].
+[18] identify order sensitivity in prompting.
+The class accuracy imbalance has multi-faceted causes, from pre-training data, such as predilections for common tokens in pretraining [19], to a liking for specific patterns or answers [20].
+Other causes include choices of prompt templates [21, 22, 23], demonstrations [24], and prompt orders [25].
+Unintended LLM behaviors may also exacerbate it, such as sycophancy[26, 27, 28].
 6 Discussion
@@
 In particular, with the same class probability vectors as input, we build a single linear adapter (LA) network with adapter dimension of 64, and train it towards the same objective as DNIP; we also follow the batch calibration method [29] to estimate a calibration term using test examples.
-Figure 8: Comparisons to more methods.
-Figure 8 shows the comparison results. fine-tuning a simple linear adapter does not work well on the class output probabilities. The proposed method outperforms both methods in accuracy and COBias.
-Using Chain-of-Thought Prompting to Improve LLM Predictions.
-Some may argue that a simple remedy for biased predictions is Chain-of-Thought (CoT) prompting, which elicit more reasoning in LLMs.
-However, Turpin et al.
-[23] points out that CoT explanations can be unfaithful, as they can be biased towards class A when prompts always put letter A as the answer, causing 36% accuracy drops on BBH [37].
-We add that prompt-based models may fall for pattern matching instead of recalling learned factual knowledge [25].
-Therefore, CoT may not help due to unfaithful explanations.
-Using Calibration Methods to Mitigate ICL Biases.
-Previous calibration methods reduce biases caused by prompting.
-For example, Zhao et al.
-[24], Zhou et al.
-[36] use content-free or content-based prompts to measure an offset and calibrate class probabilities by removing it.
-This may be good for mitigating bias towards a single class, but it does not consider the interaction between class predictions, e.g., pairwise class accuracy bias studied in this work.
-Another hazard is using prompts to compute the offset, which may cascade biases brought by a certain or several prompt templates, and biases inherited from pretraining may not be reflected.
-Instead, we correcct biased predictions without more prompting, but we differ in that our correction weights are explicitly optimized on COBias and accuracy metrics, while the cluster-based decision boundary is implicitly learned.
+Figure 6 shows the comparison results.
+fine-tuning a simple linear adapter does not work well on the class output probabilities.
+The proposed method outperforms both methods in accuracy and COBias.
 Evaluating Closed LLMs.
-While closed LLMs such as ChatGPT and GPT-4 [38] are interesting to evaluate, they do not return logits/softmax probabilities over the entire vocabulary at an output token.
-For example, ChatGPT returns log probabilities of up to 20 most likely output tokens, and may not cover the probability for every class that is needed in our class prediction debiasing.
-Optimization Algorithms for Integer Programming.
-The classic solvers for integer programming include operational research techniques such as Branch and Bound (BnB), commonly used for linear integer programming problems.
-However, it is difficult for these methods to solve our DNIP model, which contains discontinuous functions and is non-differentiable.
-In this case, a series of metaheuristic algorithms are more suitable, solversincluding Simulated Annealing (SA), Genetic Algorithms, Ant Colony Optimization, and Particle Swarm Optimization.
-These algorithms all belong to the same category, and each meta heuristic has their own pros and cons.
-Therefore, SA in this paper has a clear purpose of tackling our NIP mathematical model that classical operational research methods cannot solve.
-Our Motivation Is Different from Classical Post-hoc Corrections.
-Some may contend that achieving equitable accuracies across all classes is a well-explored issue in standard machine learning classifiers.
-However, it is crucial to see that class prediction accuracy imbalances must be understood within their specific context.
-The accuracy bias in LLMs’ outputs arises from fundamentally different causes than the unequal class accuracies seen in potentially overfitted traditional classifiers.
-In the former, biases stem from prompts and pre-training, while in the latter, the imbalance is driven by skewness in the supervised training data.
-This distinction underscores why our approach is uniquely applied to LLMs’ output token class probabilities.
+While closed LLMs such as ChatGPT [30] are interesting to debias, ChatGPT returns log probabilities of top 20 most likely tokens, which may not cover the probability for every class that is needed as inputs to DNIP.
+Appendix H has more discussions on optimization algorithms for integer programming, how our motivation is different from other classical post-hoc correction methods, using calibration methods to mitigate ICL biases, and using Chain-of-Thought prompting to improve LLM predictions.
 7 Conclusion
-This work tackles the imbalance in class prediction accuracy of language models.
-We introduce COBias and reveal that LLMs of varied scales and families exhibit large COBias at inference.
-To minimize COBias and push for higher accuracy, we view debiasing as an inference-time combinatorial optimization problem and propose DNIP, which optimizes class correction weights by jointly minimizing evaluation scores by COBias and maximizing accuracy.
-Empirical evaluations on three LLMs across seven multi-class NLP classification tasks demonstrate that DNIP simultaneously achieves lower test COBias and higher test accuracy over conventional ICL.
-We advocate that reducing COBias is a direction to gain more accurate, reliable LLM predictions.
-For future works, we may extend our studies to more tasks and other modalities.
-We aim to find simple inference-time plug-in solution for further boosting language model prediction performances.
+We introduce COBias to quantify class accuracy imbalance for language models, and emphasize the need to minimize COBias to push for more reliable predictions while maintaining good overall accuracy.
+We correct ICL output class probabilities towards this goal as a combinatorial optimization problem, proposing DNIP to optimize class correction weights.
+Evaluations on three LLMs across seven classification tasks demonstrate that DNIP achieves both lower COBias and higher overall accuracy over original ICL results.
 Appendix A Other Examples of Over-prediction and Under-prediction in LLMs
@@
 We define $H(c_{i})$ as below.
-$\textrm{odd}_{i}=\verb|odd_class|(c_{i}),$ (15)
 $H(c_{i})=\frac{\displaystyle\sum\nolimits_{m}\mathbbm{1}\{y_{m}=c_{i},\hat{y}_{m}=\textrm{odd}_{i}\}}{\displaystyle\sum\nolimits_{k}\mathbbm{1}\{y_{k}=c_{i}\}},$ (11)
@@
 We compare the effectiveness of using COBias, COBiassingle, and $H$ in the DNIP objective, and refer to the respective objectives as ZCB (Eq.
-4), ZCBs (Eq.
-14), and ZH (Eq.
-17).
-Detailed test scores for the benchmark datasets on Llama-2-13B, including class accuracies, are shown in Tables 7 to 13.
+3), ZCBs, and ZH (Eq.
+12).
+Detailed test scores for the benchmark datasets on Llama-2-13B, including class accuracies, are shown in Tables 8 to 14.
 In short, ZH still leaves some classes under-predicted, as examplified by the individual test class accuracies, and test COBias (or test COBiassingle) scores on TREC, RTE, etc.
@@
 […]
-Appendix D Additional Heatmaps for Pairwise Accuracy Differences
-Figure 9 and 10 demonstrate that more balanced class accuracies are achieved with DNIP for smaller LLMs of Llama-2-7B and GPT-2-XL.
+In addition, the brute force strategy to find optimal correction weights is enumerated search, where the number of weight combinations to explore is $K^{N}$ for an example; for $M$ examples, it results in a computational complexity of $\mathcal{O}(K^{N}\cdot M)$, which can not finish in polynomial time.
+When $M$ or $N$ is large, enumerated search is computationally intensive.
+Considering this, we did not adopt the brute force strategy.
+Appendix D Dataset Preprocessing
+We take the full training set and obtain ICL class probabilities with ground-truth labels for DNIP optimization, except for AGNews, DBpedia, and DDI, for which we randomly select 10,000 training examples, since the datasets are quite large.
+We further split the training set into 95% optimization set and 5% development set.
+Test evaluations are performed on each test set against overall accuracy and COBias, except for AGNews, for which we evaluate on 5,000 randomly selected test examples.
+Appendix E Performance Improvements vs. Number of Optimization Examples
+We evaluate DNIP with varying optimization examples on Llama-2-13B.
+We find that DNIP requires only a small optimization set of 10 examples for a fair improvement in both COBias and accuracy, and it exhibits a further, emergent COBias reduction at several thousands optimization examples.
+Figure 8: Relative test COBias reduction; higher is better.
+Test COBias reduction reaches 9% to 72% at 10 optimization examples, and continues to grow with more optimization examples.
+Figure 8 shows the relative test COBias reduction over ICL results.
+At 10 examples, DNIP can achieve a relative COBias reduction ranging from 9% to 72%.
+We observe an overall trend of COBias reduction except for DDI and PubMedQA, who have a spike at 100 examples, but DDI rises again after 500 examples while PubMedQA levels off (PubMedQA full optimization size: 950).
+An explanation is that DNIP prioritizes COBias reduction more than accuracy increase at fewer examples; at more examples, the accuracy-related objectives show more influence.
+It also indicates that domain-specific datasets are more sensitive to the trade-off.
+We further notice a surge in reductions from 1,000 to full for DDI, SST-5, and DBpedia (full optimization sizes: 9,500, 8,116, 9,500), suggesting that around 8,000 examples are needed for a further emergent reduction - PubMedQA may see further COBias reduction if more optimization examples are available.
+Test accuracy is improved with only 10 optimization examples, and plateaus at around 500 examples.
+Figure 9 shows that, at 10 examples, the relative increase in accuracy ranges from 0.2% to 11% except DDI, which is 160%.
+At 100 examples, the increase ranges from 3% to 208% (DDI, $7.2\%\rightarrow 18.7\%$).
+The increase continues to grow and stabilizes at around 500 examples.
+At 1,000 examples, test accuracy for DDI rises to 52%, which is more than 6 times of the ICL accuracy, showing that DNIP is effective for this challenging biomedical NLP task.
+Figure 9: Relative test accuracy increase; higher is better.
+Moreover, we observe for RTE a declining trend of accuracy increase, which matches its growing trend of COBias reduction, reflecting the trade-off between the two objectives.
+We also note a slight decrease in accuracy improvement for SST-5 and DDI at full size, which matches their sudden increase in COBias reduction, verifying that DNIP has emergent COBias reduction abilities with thousands of optimization examples.
+Overall, DNIP improves both accuracy and COBias at 10 examples, and achieves acceptable COBias reduction and accuracy improvement at 100 examples, with test COBias reduced by 57% and test accuracy improved.
+For detailed scores, please refer to Appendix I.
+Appendix F Additional Heatmaps for Pairwise Accuracy Differences
+Figure 10 and 11 demonstrate that more balanced class accuracies are achieved with DNIP for smaller LLMs of Llama-2-7B and GPT-2-XL.
 Figure 10: Comparisons of pairwise test class accuracy differences before and after applying DNIP on GPT-2-XL; the pinker the higher accuracy difference, the greener the lower accuracy difference; average accuracy over three runs are used.
 Figure 11: Comparisons of pairwise test class accuracy differences before and after applying DNIP on Llama-2-7B; the pinker the higher accuracy difference, the greener the lower accuracy difference; average accuracy over three runs are used.
-Appendix E Additional Ablation Analysis
-We show ablation analysis on GPT-2-XL and Llama-2-7B in Figure 11 and 12.
+Appendix G Additional Ablation Analysis
+We show ablation analysis on GPT-2-XL and Llama-2-7B in Figure 12 and 13.
 Figure 12: Ablations on objective function, GPT-2-XL, demonstrating that $z_{1}+\beta z_{2}-\tau z_{3}$ achieves a balance point between accuracy and COBias.
 Figure 13: Ablations on objective function, Llama-2-7B, demonstrating that $z_{1}+\beta z_{2}-\tau z_{3}$ achieves a balance point between accuracy and COBias.
-Appendix F Full Table with Varying Number of Optimization Examples for DNIP
-The full results for all three models across varying optimization set sizes are shown in Table 14.
+Appendix H More Discussions
+Optimization Algorithms for Integer Programming.
+The classic solvers for integer programming include operational research techniques such as Branch and Bound (BnB), commonly used for linear integer programming problems.
+However, it is difficult for these methods to solve our DNIP model, which contains discontinuous functions and is non-differentiable.
+In this case, a series of metaheuristic algorithms are more suitable, solversincluding Simulated Annealing (SA), Genetic Algorithms, Ant Colony Optimization, and Particle Swarm Optimization.
+These algorithms all belong to the same category, and each meta heuristic has their own pros and cons.
+Therefore, SA in this paper has a clear purpose of tackling our NIP mathematical model that classical operational research methods cannot solve.
+Our Motivation Is Different from Classical Post-hoc Corrections.
+Some may contend that achieving equitable accuracies across all classes is a well-explored issue in standard machine learning classifiers.
+However, it is crucial to see that class prediction accuracy imbalances must be understood within their specific context.
+The accuracy bias in LLMs’ outputs arises from fundamentally different causes than the unequal class accuracies seen in potentially overfitted traditional classifiers.
+In the former, biases stem from prompts and pre-training, while in the latter, the imbalance is driven by skewness in the supervised training data.
+This distinction underscores why our approach is uniquely applied to LLMs’ output token class probabilities.
+Using Calibration Methods to Mitigate ICL Biases.
+Previous calibration methods reduce biases caused by prompting.
+For example, Zhao et al.
+[19], Zhou et al.
+[29] use content-free or content-based prompts to measure an offset and calibrate class probabilities by removing it.
+This may be good for mitigating bias towards a single class, but it does not consider the interaction between class predictions, e.g., pairwise class accuracy bias studied in this work.
+Another hazard is using prompts to compute the offset, which may cascade biases brought by a certain or several prompt templates, and biases inherited from pretraining may not be reflected.
+Instead, we correcct biased predictions without more prompting, but we differ in that our correction weights are explicitly optimized on COBias and accuracy metrics, while the cluster-based decision boundary is implicitly learned.
+Using Chain-of-Thought Prompting to Improve LLM Predictions.
+Some may argue that a simple remedy for biased predictions is Chain-of-Thought (CoT) prompting, which elicit more reasoning in LLMs.
+However, Turpin et al.
+[25] points out that CoT explanations can be unfaithful, as they can be biased towards class A when prompts always put letter A as the answer, causing 36% accuracy drops on BBH [32].
+We add that prompt-based models may fall for pattern matching instead of recalling learned factual knowledge [20].
+Therefore, CoT may not help due to unfaithful explanations.
+Appendix I Full Table with Varying Number of Optimization Examples for DNIP
+As an addition to Appendix E, the full results for all three models across varying optimization set sizes are shown in Table 15.
 Opt. set Eval. Metric AGNews DBpedia SST-5 TREC RTE DDI PubMedaQA